### v1

# MAD-Sherlock: Multi-Agent Debates for Out-of-Context Misinformation Detection

Abstract
One of the most challenging forms of misinformation involves the out-of-context (OOC) use of images paired with misleading text, creating false narratives. Existing AI-driven detection systems lack explainability and require expensive finetuning. We address these issues with MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection. MAD-Sherlock introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making. Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning. Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that MAD-Sherlock boosts performance for both experts and non-experts. These results position MAD-Sherlock as a powerful tool for autonomous and citizen intelligence applications.
1 Introduction
Our growing dependence on online channels for news and social networking has been complemented by a surge in exploits of digital misinformation (Aslett et al., 2024; Hasher et al., 1977; Brashier & Marsh, 2020). While many manipulation techniques pose serious threats, one of the most prevalent methods for creating fake online content is the out-of-context (OOC) use of images (pbs, ). This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise. Indeed, the problem of OOC misinformation detection requires a complex understanding of the relationship between the text and image and the ability to identify when they do not go together. Identifying these minute inconsistencies is a time-consuming and high-effort task for humans. A study by Sultan et al. (2022) shows that time pressure reduces the ability of human beings to detect misinformation effectively, further adding to the scalability issues in human expert detection.
Therefore, attention has turned to AI-driven tools that can help human experts recognise instances of OOC image-based misinformation at scale. Unfortunately, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), which target detecting manipulations such as PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated (or manipulated) fake images called Deepfakes (mit, ), rely on spotting artifacts from image or text tampering. In contrast, OOC detection demands cross-contextual reasoning, as the deception arises from the misalignment between the legitimate image and its falsely associated textual content.
Pretrained Large Multimodal Models  (Liu et al., 2024b; OpenAI & et al., 2024; Li et al., 2019; Radford et al., 2021, LMMs) provide a promising direction for detecting OOC use of images for their ability to process both text and image content in tandem. However, using LMMs directly for OOC detection presents several challenges, particularly in the news domain. For instance, news articles often include images that are not directly related to the article’s content. An article about the 2024 U.S. presidential candidates, for example, might feature a close-up of Donald Trump from an unrelated online database. Although the image was taken outside the election period, it is not considered OOC since it doesn’t misrepresent the article’s context. Such cases complicate LMMs’ ability to accurately identify OOC usage based solely on their pre-trained knowledge, as this knowledge may be outdated or insufficiently detailed.
Figure 1: Overview of MAD-Sherlock: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not. After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
Moreover, even with recent advancements, LMMs are capable of hallucinating and, hence, generating false information (Bai et al., 2024; Liu et al., 2024a). While rapidly improving, they sometimes fail to understand user instructions and intent correctly. We show that off-the-shelf LMMs indeed suffer from these issues, reducing their ability to detect OOC misinformation in practice. While prior work (Qi et al., 2024) has shown that off-the-shelf models can be improved using task-specific fine-tuning, this approach is resource-intensive and requires continual updating to keep up with recent events. Moreover, detecting OOC images only solves part of the problem. The real value lies in being able to explain the OOC use of pictures in human-readable form. It can be instrumental for human validators to observe the model’s line of logic and gain better insight into, and trust in, the classification process.
In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3). Specifically, our framework MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allow for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b). In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of MAD-Sherlock as an interactive tool for human experts. To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
We perform a comprehensive empirical evaluation of our method (see Section 4), including the study of multiple debate configurations. To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance. We find that MAD-Sherlock outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies. We identify both access to external information retrieval and complete freedom of opinion as key ingredients to MAD-Sherlock’s performance. Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
2 Related Work
Recent work has focused on using joint image-text representations to classify an instance as OOC. Aneja et al. (2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar. They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption. During inference, they use the semantic similarity between the two captions to classify them as OOC or not. The increased reliance on textual content limits the capabilities of this approach. This work also does not provide explanations for model predictions and is, therefore not interpretable. Moreover, this method works for image caption pairs where captions have information about objects in the image. This is not always the case with news articles (our domain of application), where captions can often just be related to the main content of an article rather than precisely describing the objects in the image. Appendix  A.2 shows an example of the same.
Abdelnabi et al. (2022) present the Consistency Checking Network (CCN) in which they emulate different aspects of human reasoning across modalities for misinformation detection. This method uses evidence related to the image-text pair aggregated from the Internet. The CCN consists of memory networks to assess the consistency of the image-caption pair against the retrieved evidence and a CLIP (Radford et al. (2021)) component to evaluate the consistency between the image and caption pair. The use of external evidence to better inform model decisions is an important idea and also explains the superior classification performance of CCN when compared to other methods. This method also lacks the explainability component.
Zhang et al. (2024) extend the neural symbolic method (Yi et al. (2019); Zhu et al. (2022)) to propose an interpretable cross-modal misinformation detection model to provide supporting evidence for the output prediction. They use symbolic graphs based on the Abstract Meaning Representation (Banarescu et al. (2013)) of textual and visual information to detect OOC image use. Zhou et al. (2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks are used to learn features of text and visual news representations. Their representations and relationships are jointly learned and used to predict fake news. Wang et al. (2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated. EANN uses adversarial training to learn multi-modal features independent of news events. These methods require pretraining from scratch and, therefore, don’t benefit from the advanced reasoning capabilities and world knowledge of large pretrained models.
Shalabi et al. (2023) use synthetic multi-modal data to establish the authenticity of image-text pairs. They use BLIP-2 (Li et al. (2023b)) to generate a caption for the original image and Stable Diffusion (Rombach et al. (2022)) to generate an image for the given original caption. This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text. This method relies on synthetic multi-modal data generation, which not only adds an additional computational overhead but also increases dependence on often unreliable synthetically generated data. Therefore, this method can suffer from issues related to generation models, including potential biases that these models may possess. This method also lacks interpretability.
Sniffer (Qi et al. (2024)) is the closest to our work. It uses the InstructBLIP (Dai et al. (2023)) model to detect OOC image use and provide an explanation for its prediction. It makes use of internal and external knowledge using entity extraction APIs and image-based web searches. Information from all the sources is given to an LLM to predict and explain if an image has been used OOC. Sniffer only uses basic textual information, such as news article titles from websites, to form its external knowledge base. It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
Lin et al. (2024) use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful. These explanations are then used to train a small language model as a judge to determine whether the image and text that make up the meme are actually harmful. This work does not allow agents to have flexibility of opinion. There are always two agents, and each one is provided a stance to defend. Moreover, a judge decides the final outcome of the debate and needs to be trained on data from the debate. This method also does not benefit from external retrieval, and therefore, the debating agents are not aware of the crucial external context related to the input. Finally, this work is related to harmful meme detection and does not concern the problem of misinformation detection in the news domain, which likely requires more intricate contextual analysis, including of external context.
3 Methodology
We present an explainable misinformation detection system, MAD-Sherlock, which jointly predicts and explains instances of misinformation. Figure 1 illustrates our approach. To the best of our knowledge, all prior work except Qi et al. (2024) provide predictions without explanations, and no prior work uses multiple models to approach this problem. We present a novel methodology that involves multiple multi-modal models debating against each other in order to decide if an image-text pair is misinformation or not. In this work, we aim to answer the question:
Can debating multi-modal models, when equipped with external context, be used to solve the problem of explainable misinformation detection by picking up on minute contextual inconsistencies?
We use detailed external information to inform the model’s predictions through the external retrieval module, which utilises reverse image-based search in order to provide the agents with external real-world context related to the image-text pair. We carry out our experiments with the GPT-4o (OpenAI ) model to achieve state-of-the-art performance on the misinformation detection task while also providing detailed and coherent explanations for the predictions. We achieve this without any domain-specific fine-tuning, thus ensuring easier and faster generalization to other domains in addition to low computational overhead.
3.1 Debate Modelling
Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways. We explore multiple debating strategies to structure the conversation between agents. Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s). We observe this facilitates more involved and detailed discussions among the models.
Figure 2: Debating Strategies: We experiment with multiple debating strategies. The asynchronous debate setup where agents argue one after the other and take turns presenting their arguments is the best configuration.
Asynchronous Debate (not) against Human:
We define an asynchronous debating strategy in which models wait for the other participants’ responses before generating their own. Figure 2 (a) and (b) show synchronous and asynchronous debating structures, respectively. While synchronous debates, where all participants speak at once, can be faster and computationally more efficient, we opt for an asynchronous setting where each model response is based on previous responses of other models. We observe that this method of structuring the debate works better since models are able to pick up on contextual ambiguities in their responses in a more organised and structured way which is crucial to the process of misinformation detection. An important point to note for this setup is the way we structure model prompts. The debating models are not aware that they are debating other AI agents. The prompts are structured such that each debating model believes it is talking to a human.
Judged Debate:
We also experiment with an asynchronous debate setup with a judge. Figure 2 (c) shows this setup. In this setup, models participate in an asynchronous debate as usual however, the final decision is made by a judge at the end of the debate. Models are incentivised to structure their arguments in a way that makes them most convincing to the judge. We structure this debate configuration similar to Khan et al. (2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
Actor-Skeptic:
In this setup, only one agent, the actor, is tasked with deciding whether a given image-text pair is misinformation. The agent generates a response which a skeptic then evaluates. The skeptic is tasked with finding logical errors in the actor’s argument and asking follow-up questions to disambiguate the actor’s response. It is important to note that neither the skeptic nor the actor has access to the ground truth. This setup does not benefit from an ensemble since both models assume different roles and only one agent is tasked with generating the final answer.
Debate with Disambiguation:
Improving on the actor-skeptic method, in this setup, we allow all agents to act as actors and skeptics. Models are tasked with not only generating their own responses but also disambiguation queries to refine further or refute the other agents’ responses. These disambiguation queries are then used to search the Internet to obtain information to refine model outputs further. We are the first to propose this debate setup and while it does not achieve the best results in this work, we believe future research can greatly benefit by refining this setup further.
We empirically identify the best debate set up to be the asynchronous debate where the model believes it is debating against a human and not an AI agent.
3.2 Prompt Engineering
The debate structure is substantiated through prompt engineering. Figure 1 shows that the first stage of our method requires for each AI agent to generate an independent response to whether the given image-text pair is misinformation. Each agent must take into account the external context related to the image obtained through the external information retrieval module. Specifications of the various prompts used in this work can be found in Appendix A.3. An initial prompt provides the agent with a summary of the news articles related to the image and, based on it, asks the agent to classify the image-text pair as misinformation or not. The prompt asks the agent to focus on certain details in the image, such as watermarks, flags, etc. We observe that images used in news articles often contain minute yet crucial details which can be used to inform the final decision about whether the image actually belongs to the news articles. Therefore, prompting the agent to pay special attention to these details further helps detect inconsistencies.
Once the agents have formed independent opinions about the image-text pair, they must then participate in a debate. While the same prompt can be used for each debate round, we provide a different prompt for the first round to allow each agent to understand the changing nature of the conversation. Responses from other agents are provided as a part of the prompt, and the agent is asked to agree or disagree. The agent must also clearly state the reasoning process behind its argument. The prompt further requires the agent to identify ambiguities in the other agents’ reasoning. This allows the agent to closely analyse the responses from other agents and use them to inform its own decision.
A separate prompt is then used to facilitate all rounds of the debate following round one. It takes into account the other agent’s suggestions from the previous round and, after refining its own response, asks the agent to point out any inconsistencies in this new response. We note that in such a scenario, agents are prone to simply agreeing with each other and repeating the other agent’s response. This is especially the case when agents believe they are conversing with a human. However, this tendency of the agents to easily give up their own opinions and simply agree with the other participants does not facilitate an advantageous debate, and the agents do not discover any new information related to the image-text pair. Therefore, the debate prompt also contains clear and explicit instructions asking the agent not to simply agree with the presented response unless it has an acceptable reason to do so. We find this helps the agents develop stronger stances and reluctance to blind agreement.
3.3 External Information Retrieval
Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al. (2023)). This summary is then passed to the debating agents as a part of the initial prompt.
Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame). Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al. (2022)) to supplement external information related to an image-caption pair. However, we find this information lacking in detail since it is limited to the title of a news article. Agents can greatly benefit from the knowledge of the entire news article and its content rather than just the title when making a decision about whether a given image-caption pair, when considered in the context of the news article, is misinformation. To this end, we propose our own external information retrieval module. We observe a significant improvement in accuracy after incorporation of the external information retrieval module into the pipeline. The module is implemented in two stages:
3.3.1 API-Based Information Retrieval
The Bing Visual Search API (vis ) is used for the task of obtaining web pages related to a given image. A given image from the dataset is used to obtain a list of web pages completely and partially related to the image. We take the top three matching web pages in which the image appears. We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used. Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al. (2021)), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results. In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base. Since this is not the case for a significant fraction of examples in the dataset, it does not adversely affect system performance.
3.3.2 Summarization using LLM
Once the top three web pages have been identified, we scrape the text from the web pages to obtain the textual information related to the context in which the image appears on the Internet. The compiled textual information is often too long to be passed directly to the agent, and hence, we use the Llama-13B (Touvron et al. (2023)) language model to summarize this information. The summaries obtained from the LLM only focus on the most important parts of the text and hence also allow agents to develop a more focused understanding of the external context. While this method of summarization works for most samples in our dataset, there are some examples where the obtained web pages are not in the English language, and the LLM struggles with summarization. In this regard, we add an additional check that ignores text from web pages in languages other than English. While this restricts our system to the English language, it does not adversely affect system performance due to the distribution of the dataset, which consists of images mostly taken from English-language news articles. Multi-lingual support can be achieved by first translating text to English and then summarizing it.
3.4 Coherent Reasoning
All the different components of MAD-Sherlock are brought together in this stage of the pipeline. Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same. The agents also have access to external information related to the image through the external retrieval module. The final decision of the system is obtained once the debate terminates, which is after a certain number of debate rounds or after all agents converge to a common response, whichever is earlier.
4 Experiments and Results
4.1 Dataset
We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al. (2021)). The dataset is built based on the VisualNews (Liu et al. (2020)) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian and The Washington Post. The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair. CLIP (Radford et al. (2021)) is used to retrieve semantically similar images for a given caption. We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples. The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively.
4.2 Experimental Setup
All experiments were run on 8 A40 (46GB) Nvidia GPU server.
Debate Setup:
We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al. (2024b)). The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
External Retrieval Module:
We use the Bing Visual Search API (vis ) to run an image-based reverse search. Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new ). Finally, we use Llama-13B (Touvron et al. (2023)) to summarise the text obtained from the top $k=3$ web pages. This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
Baselines and Prior Work:
We compare MAD-Sherlock to existing pretrained multi-modal baselines including CLIP (Radford et al. (2021)), VisualBERT (Li et al. (2019)), InstructBLIP (Dai et al. (2023)) and LLaVA (Liu et al. (2024b)). We also compare performance against GPT-4o (OpenAI & et al. (2024); OpenAI ). The models are presented with the image and caption pair and asked if the pair is misinformation. The models are further prompted to explain their reasoning. We also show results for two baseline methods trained from scratch, namely EANN (Wang et al. (2018)) and SAFE (Zhou et al. (2020)). We further compare MAD-Sherlock to DT-Transformer (Papadopoulos et al. (2023)), CCN (Abdelnabi et al. (2022)), Sniffer (Qi et al. (2024)), VINVL (Huang et al. (2024)), SSDL (Mu et al. (2023)) and Neuro-Sym (Zhu et al. (2022)).
4.3 Results
We present results for the experiments conducted to select the best debate setup as well compare the performance of MAD-Sherlock against existing methods. We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
4.3.1 Comparing Debate Setups
We compare multiple debating setups using the LLaVA model, to select the best one for comparison with other works and further experimentation.
Debate Setup Accuracy Precision Recall
Async_DebateAI (believes debating AI) 75.2 54.5 86.4
Async_Debatehuman (w/o external info) 77.1 68.4 89.3
Async_Debatehuman (w external info) 86.2 82.6 90.6
Actor-Skeptic 69.5 66.1 69.4
Judged Debate 66.7 66.7 61.5
Debate with Disambiguation 77.8 74.7 82.6
Table 1: Performance comparison between different debate setups: The Async_Debatehuman where the model has external context and believes it is debating a human being is the best setup.
Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations. We also report results for the Asynchronous Debate setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain. The external retrieval of information significantly boosts performance. All following debate set-ups, therefore, use external information related to the image-caption pair as a part of their initial prompt. We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent. Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human. Further, the Asynchronous Debate setup benefits from the ensemble of agents which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses. The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought. We believe this accounts for the counter-intuitively performance of this method where agents perform worse with more information. The judged debate setup focuses on enforcing agents to structure their responses in a way that will convince a judge. The agents also debate with opposite stances and do not have the option of changing their stance mid-debate. This can further confuse the judge and lead to incorrect decisions. This is resolved in the Async_Debatehuman set up where agents are given complete freedom over their initial opinions, as well as their opinions during the debate. If they believe they are convinced by the other agents’ arguments, they can choose to change their response and the debate ends. Based on the results from Table 1, we choose the best-performing debate set-up, i.e. Async_Debatehuman (with external information) as the debate configuration for further experimentation and comparison.
4.3.2 Performance Comparison
We present our results on the NewsCLIPpings dataset against existing out-of-context detection methods discussed in section 4.2.
Model Accuracy
SAFE 50.7
EANN 58.1
VisualBERT 54.8
CLIP 62.6
InstructBLIP 48.6
LLaVA 57.1
GPT-4o 70.7
DT-Transformer 77.1
CCN 84.7
SSDL 65.6
VINVL 65.4
Neuro-Sym 68.2
Sniffer (w finetuning) 88.4
Sniffer (w/o finetuning) 84.5
MAD-Sherlock (ours) 90.17
Table 2: Performance comparison between our model and baselines: MAD-Sherlock (with GPT-4o) out performs all related work.
Table 2 shows the comparison between our system and existing methods. We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al. (2024); OpenAI ) model. Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset. While we do not provide a quantitative assessment of explanations by MAD-Sherlock, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines. We leave the detailed analysis of these explanations and the development of the associated metrics as future work.
We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work. This can be attributed to the necessity for external context for misinformation detection in the news domain and the lack of diverse perspectives that arise naturally in a multi-agent framework. Therefore, these standalone models, while promising, currently are unable to detect misinformation effectively. These models require additional integration into more comprehensive pipelines, as done in this work. In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al. (2020)) and EANN (Wang et al. (2018)) perform worse than pretrained multi-modal models. This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
5 User Study
We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation. While it is easy to quantify model performance in terms of misinformation detection, there are no effective metrics to assess the quality of the explanations generated by the model. Therefore, in order to perform a thorough analysis of the system performance, a user study is essential. For a deeper analysis we further grouped the participants based on their profession into three groups, namely: Journalists, AI Academics (studying AI) and Others. Further details regarding the study setup and participant groups can be found in Appendix A.4.
In the study, participants were shown ten image-text pairs and were asked to decide if the image and caption, when considered together, were misinformation or not. They were also asked to provide a confidence rating for their answer on a scale of 0-10, with 10 being the highest confidence level. For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair. These AI insights were the final output explanations from MAD-Sherlock. Participants were then asked to reconsider their answers in light of the new information from the AI agent. Table 3 shows that average system performance is better than the average human performance for both cases where the participants have access to AI insights and where they do not. Therefore, MAD-Sherlock can be used as a reliable assistive tool for OSINT research for detecting and explaining misinformation with little or no human intervention.
Study Setup Average Accuracy
Humans $60.3\pm 13.5$
Humans+MAD-Sherlock $76.7\pm 12.2$
MAD-Sherlock $\mathbf{80.0\pm 0.0}$
Table 3: Performance comparison between different study setups: MAD-Sherlock outperforms humans with and without AI assistance.
We further observe that the average human accuracy for the misinformation detection task increases by more than 27% concretizing the fact that AI insights from our model do actually improve human efficiency in detecting misinformation. We also observe interesting patterns for group-wise analysis which we believe would be valuable for future work. Table 4 shows that the performance of all groups improves significantly and is not far off from that of professional journalists. The average confidence level (out of 10) is comparable across all the groups before and after considering MAD-Sherlock insights and generally increases. Therefore, we conclude that MAD-Sherlock can significantly uplift non-expert performance and hence can be useful in citizen intelligence applications.
Avg acc $\uparrow$ Avg conf $\uparrow$ Avg acc $\uparrow$ Avg conf $\uparrow$
Group (only human) (only human) (with (with
MAD-Sherlock) MAD-Sherlock)
Journalists $70.0\pm 1.4$ $4.3\pm 2.1$ $82.2\pm 0.9$ $5.3\pm 1.3$
AI Academics $60.7\pm 1.4$ $3.2\pm 0.8$ $79.3\pm 1.3$ $5.8\pm 1.4$
Others $56.7\pm 1.5$ $3.9\pm 1.2$ $71.7\pm 1.1$ $5.8\pm 1.4$
Table 4: Performance comparison between different participant groups: All groups show performance improvement with MAD-Sherlock. AI Academics are able to perform nearly at par with professional journalists after considering insights from MAD-Sherlock.
6 Conclusion and Future Work
Misinformation detection has become a pressing issue in recent times. With the ever-advancing capabilities of vision and language models, the detection of OOC image use has become a very difficult task. In this work, we explore the question of whether it is possible for multiple AI agents to pool their contextual knowledge and converge to a common prediction in order to identify instances of misinformation. We identify Asynchronous_Debatehuman as the most optimal communication setup for AI models. We observe significant performance improvement when the models believe they are debating against a human instead of another AI agent. We observe that in this setup, models tend to be more involved and open to changing their opinions. Our method also allows for agents to have freedom of opinion which they may change mid-debate. Agents in such a setting show enhanced abilities to critically evaluate an argument and pick up on minute inconsistencies.
Our final system, MAD-Sherlock, achieves state-of-the-art performance on the misinformation detection task. Further, owing to our advanced external retrieval module, MAD-Sherlock provides clear, coherent and detailed explanations. As a result, MAD-Sherlock significantly improves the OOC misinformation detection performance of both human experts, and non-experts.
We identify several promising avenues for future research in this field. Notably, improving how agents handle disambiguation queries could significantly enhance detection accuracy and explainability by providing access to additional information. The research community would benefit from a continuously updated benchmark dataset, incorporating more recent news articles and subtler inconsistencies. A direct extension of this work involves applying our methods to video-text pairs. Since MAD-Sherlock’s explanations are currently limited to text, future research should explore multi-modal and mixed-modal explanation generation. Additionally, further experiments with different hyper-parameters, such as varying the number of agents and debate rounds, could refine the effectiveness of our approach.
Finally, while we conducted extensive user studies with MAD-Sherlock, deploying it on a larger scale in professional environments and within the citizen intelligence community will provide valuable insights into its real-world performance, uncovering new opportunities for improvement. For an analysis of limitations, please refer to Appendix A.1.
7 Acknowledgments
A part of this work was supported by UKRI DECaDE EP/T022485/1 and the UKRI grant: Turing AI Fellowship EP/W002981/1, Armasuisse Science+Technology. We would also like to thank the Royal Academy of Engineering and FiveAI.
Appendix A Appendix
A.1 Limitations
Despite the strong performance of MAD-Sherlock, several limitations remain. First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve. Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration. Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
Additionally, while our user studies provided valuable insights, large-scale deployment in diverse, real-world settings, such as professional or citizen intelligence environments, is necessary to fully assess the method’s robustness and scalability. Finally, our dataset, though comprehensive, primarily focuses on English-language news, limiting the generalizability of the system across non-English contexts.
Another important limitation is the potential risk that open-sourcing MAD-Sherlock might allow adversaries to train models specifically designed to counter or evade detection by our system. As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness. However, despite these risks, we believe that open-sourcing remains the right path forward. Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining MAD-Sherlock’s effectiveness in the long term. The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of MAD-Sherlock.
A.2 Sample Image-Caption Pair in the News Domain
Figure 4: Russian President Vladimir Putin has called Ukraine’s move into Kursk a “major provocation”. Image and caption taken from the BBC article here (Accessed at 17:43 on Aug 11, 2024): https://www.bbc.co.uk/news/articles/cze5pkg5jwlo
A.3 Prompts for MAD-Sherlock
This is a summary of news articles related to the image: {}
Based on this, you need to decide if the caption given below belongs to the image or if it is being used to spread false information to mislead people.
CAPTION: {}
Note that the image is real. It has not been digitally altered.
Carefully examine the image for any known entities, people, watermarks, dates, landmarks, flags, text, logos and other details which could give you important information to better explain your answer.
The goal is to correctly identify if this image caption pair is misinformation or not and to explain your answer in detail.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 5: Initial prompt for independent opinion formation and response generation
This is what I think: {}.
Do you agree with me?
If you think I am wrong then convince me why you are correct.
Clearly state your reasoning and tell me if I am missing out on some important information or am making some logical error.
Do not describe the image.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 6: Prompt for Debate Round 1
I see what you mean and this is what I think: {}.
Do you agree with me?
If not then point out the inconsistencies in my argument (e.g. location, time or person related logical confusion) and explain why you are correct.
If you disagree with me then clearly state why and what information I am overlooking.
Find disambiguation in my answer if any and ask questions to resolve them.
I want you to help me improve my argument and explanation.
Don’t give up your original opinion without clear reasons, DO NOT simply agree with me without proper reasoning.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 7: Prompt for Debate after Round 1
A.4 User Study
We conduct a user study to assess the effectiveness of our model in detecting and explaining misinformation. Through this study, we aim to assess the persuasiveness of our system.
A.4.1 Setup
The user study was designed to evaluate the effectiveness of our system in detecting and explaining misinformation. While it is easy to quantify model performance in terms of misinformation detection, there are no effective metrics to assess the quality of the explanations generated by the model. Therefore, in order to perform a thorough analysis of the system performance, a user study is essential.
A total of 30 participants volunteered to participate in this study. The group of individuals included journalists from BBC as well as students and professors from the University of Oxford. Participation was completely voluntary and no personal information was used for the purpose of analysis in this study. For a deeper analysis we further grouped the participants based on their profession into three groups, namely: Journalists, AI Academics and Others. The ‘others’ category included anyone who did not belong to the first two groups. The study was conducted through a Microsoft Form. Participants were shown 10 image-text pairs and were asked to decide if the image and caption when considered together was misinformation or not. They were also asked to provide a confidence rating for their answer on a scale of 0-10, with 10 being the highest confidence level. For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair. These AI insights were the final outputs from MAD-Sherlock. Participants were then asked to reconsider their answer and again decide if the image-text pair was misinformation or not, in light of the new information from the AI agent. Participants were also required to re-evaluate their confidence score in this new answer. While it is not entirely avoidable, we did ask participants to keep aside their personal opinions of AI and consider all AI insights objectively. Participants were not allowed to access the Internet. This was done to ensure an unbiased estimate of average human performance.
The image-text pairs to include in the study were taken from the NewsCLIPpings (Luo et al. (2021)) dataset. AI insights were taken from our best-performing setup involving the GPT-4o model. Of the 10 image-text pairs presented to the participants in the study, there were 5 instances of misinformation and 5 instances of true information. Further, all model insights were true except two of them. Therefore the model accuracy for the task was 80% and we use this as the baseline accuracy to compare human performance against.
We analyse two special cases, where MAD-Sherlock argues for the wrong answer. We include these results in order to observe how persuasive our system can be even when it is wrong. We note in the instance where the image-text pair was actually misinformation and the model argued that it was not, 6 participants changed their correct responses to those suggested by MAD-Sherlock. Although this is only 5% of the participants, it still gives a significant insight into how persuasive the model can appear even when it is wrong. While the case of false negatives is important, false positives are an even more concerning matter for our problem statement. In the case where MAD-Sherlock declared the given image-text pair to be misinformation when it was not, is important to analyse. In this setting 50% of the total participants changed their answer to the wrong one, therefore believing a piece of true information to be false. In some cases where participants chose the wrong response to begin with, their confidence in the response further increased after considering insights from the system. Finally, 4 participants did not change their answer to the wrong one after considering AI insights but their confidence in their response decreased.
The average time taken to complete the study was 12 minutes and 57 seconds. The average participant was therefore able to go through 10 image-text pairs and decide if they were misinformation or not in under 13 minutes. The same task without AI insights would require extensive analysis and we project it would take between 30-45 minutes to decide if 10 image-text pairs were misinformation.
Acknowledgements
Part of this work was supported by UKRI DECaDE EP/T022485/1

### v2

@@
-MAD-Sherlock: Multi-Agent Debates for Out-of-Context Misinformation Detection
+LLM-Consensus: Multi-Agent Debate for Visual Misinformation Detection
 Abstract
@@
 Existing AI-driven detection systems lack explainability and require expensive finetuning.
-We address these issues with MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection.
-MAD-Sherlock introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making.
+We address these issues with LLM-Consensus, a multi-agent debate system for OOC misinformation detection.
+LLM-Consensus introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making.
 Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning.
-Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that MAD-Sherlock boosts performance for both experts and non-experts.
-These results position MAD-Sherlock as a powerful tool for autonomous and citizen intelligence applications.
+Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that LLM-Consensus boosts performance for both experts and non-experts.
+These results position LLM-Consensus as a powerful tool for autonomous and citizen intelligence applications.
+misinformation detection,out-of-context image use,LLMs,multimodal models,multi-agent debate,safety
 1 Introduction
@@
 This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise.
-Indeed, the problem of OOC misinformation detection requires a complex understanding of the relationship between the text and image and the ability to identify when they do not go together.
-Identifying these minute inconsistencies is a time-consuming and high-effort task for humans.
-A study by Sultan et al.
-(2022) shows that time pressure reduces the ability of human beings to detect misinformation effectively, further adding to the scalability issues in human expert detection.
-Therefore, attention has turned to AI-driven tools that can help human experts recognise instances of OOC image-based misinformation at scale.
-Unfortunately, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), which target detecting manipulations such as PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated (or manipulated) fake images called Deepfakes (mit, ), rely on spotting artifacts from image or text tampering.
-In contrast, OOC detection demands cross-contextual reasoning, as the deception arises from the misalignment between the legitimate image and its falsely associated textual content.
+OOC misinformation detection requires a nuanced understanding of the relationship between text and images and the ability to identify when they are misaligned.
+Detecting these subtle inconsistencies is time-consuming for humans, and (Sultan et al., 2022) shows that time pressure further reduces detection accuracy, exacerbating scalability issues.
+As a result, AI-driven tools have gained attention for scaling detection efforts.
+However, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), designed to detect manipulations like PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated Deepfakes (mit, ), focus on spotting artifacts from tampering.
+In contrast, OOC detection requires cross-contextual reasoning, as the deception stems from the misalignment of legitimate images with misleading text.
+Figure 1: Overview of LLM-Consensus: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not.
+After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
 Pretrained Large Multimodal Models  (Liu et al., 2024b; OpenAI & et al., 2024; Li et al., 2019; Radford et al., 2021, LMMs) provide a promising direction for detecting OOC use of images for their ability to process both text and image content in tandem.
@@
 Such cases complicate LMMs’ ability to accurately identify OOC usage based solely on their pre-trained knowledge, as this knowledge may be outdated or insufficiently detailed.
-Figure 1: Overview of MAD-Sherlock: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not.
-After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
 Moreover, even with recent advancements, LMMs are capable of hallucinating and, hence, generating false information (Bai et al., 2024; Liu et al., 2024a).
@@
 In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3).
-Specifically, our framework MAD-Sherlock: a Multi-Agent Debate system for OOC Misinformation Detection frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
-Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allow for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b).
-In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of MAD-Sherlock as an interactive tool for human experts.
+Specifically, our framework LLM-Consensus frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
+Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allows for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b).
+In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of LLM-Consensus as an interactive tool for human experts.
 To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
@@
 To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance.
-We find that MAD-Sherlock outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies.
-We identify both access to external information retrieval and complete freedom of opinion as key ingredients to MAD-Sherlock’s performance.
+We find that LLM-Consensus outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies.
+We identify both access to external information retrieval and complete freedom of opinion as key ingredients to the performance of LLM-Consensus.
 Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
@@
 Recent work has focused on using joint image-text representations to classify an instance as OOC.
-Aneja et al.
-(2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar.
+(Aneja et al., 2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar.
 They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption.
@@
 Appendix  A.2 shows an example of the same.
-Abdelnabi et al.
-(2022) present the Consistency Checking Network (CCN) in which they emulate different aspects of human reasoning across modalities for misinformation detection.
-This method uses evidence related to the image-text pair aggregated from the Internet.
-The CCN consists of memory networks to assess the consistency of the image-caption pair against the retrieved evidence and a CLIP (Radford et al.
-(2021)) component to evaluate the consistency between the image and caption pair.
-The use of external evidence to better inform model decisions is an important idea and also explains the superior classification performance of CCN when compared to other methods.
-This method also lacks the explainability component.
-Zhang et al.
-(2024) extend the neural symbolic method (Yi et al.
-(2019); Zhu et al.
-(2022)) to propose an interpretable cross-modal misinformation detection model to provide supporting evidence for the output prediction.
-They use symbolic graphs based on the Abstract Meaning Representation (Banarescu et al.
-(2013)) of textual and visual information to detect OOC image use.
-Zhou et al.
-(2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks are used to learn features of text and visual news representations.
-Their representations and relationships are jointly learned and used to predict fake news.
-Wang et al.
-(2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated.
+(Abdelnabi et al., 2022) introduce the Consistency Checking Network (CCN), which models aspects of human reasoning across modalities for misinformation detection using external evidence aggregated from the Internet.
+The CCN employs memory networks to evaluate the consistency of image-caption pairs against retrieved evidence and a CLIP component (Radford et al., 2021) to assess the alignment between the image and caption.
+The use of external evidence enhances CCN’s classification performance compared to other methods.
+However, it lacks an explainability component, providing only binary decisions without further explanation.
+(Zhang et al., 2024) extend the neural symbolic method (Yi et al., 2019; Zhu et al., 2022) to develop an interpretable cross-modal misinformation detection model that provides supporting evidence for its predictions.
+They use symbolic graphs based on Abstract Meaning Representation (Banarescu et al., 2013) to detect out-of-context image use.
+Similarly, (Zhou et al., 2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks jointly learn and analyze features and relationships between text and visual news representations to predict fake news.
+(Wang et al., 2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated.
 EANN uses adversarial training to learn multi-modal features independent of news events.
@@
 (2023) use synthetic multi-modal data to establish the authenticity of image-text pairs.
-They use BLIP-2 (Li et al.
-(2023b)) to generate a caption for the original image and Stable Diffusion (Rombach et al.
-(2022)) to generate an image for the given original caption.
+They use BLIP-2 (Li et al., 2023b) to generate a caption for the original image and Stable Diffusion (Rombach et al., 2022) to generate an image for the given original caption.
 This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text.
@@
 This method also lacks interpretability.
-Sniffer (Qi et al.
-(2024)) is the closest to our work.
-It uses the InstructBLIP (Dai et al.
-(2023)) model to detect OOC image use and provide an explanation for its prediction.
+Sniffer (Qi et al., 2024) is the closest to our work.
+It uses the InstructBLIP (Dai et al., 2023) model to detect OOC image use and provide an explanation for its prediction.
 It makes use of internal and external knowledge using entity extraction APIs and image-based web searches.
@@
 It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
-Lin et al.
-(2024) use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful.
-These explanations are then used to train a small language model as a judge to determine whether the image and text that make up the meme are actually harmful.
-This work does not allow agents to have flexibility of opinion.
-There are always two agents, and each one is provided a stance to defend.
-Moreover, a judge decides the final outcome of the debate and needs to be trained on data from the debate.
-This method also does not benefit from external retrieval, and therefore, the debating agents are not aware of the crucial external context related to the input.
-Finally, this work is related to harmful meme detection and does not concern the problem of misinformation detection in the news domain, which likely requires more intricate contextual analysis, including of external context.
 3 Methodology
-We present an explainable misinformation detection system, MAD-Sherlock, which jointly predicts and explains instances of misinformation.
-Figure 1 illustrates our approach.
-To the best of our knowledge, all prior work except Qi et al.
-(2024) provide predictions without explanations, and no prior work uses multiple models to approach this problem.
-We present a novel methodology that involves multiple multi-modal models debating against each other in order to decide if an image-text pair is misinformation or not.
-In this work, we aim to answer the question:
-Can debating multi-modal models, when equipped with external context, be used to solve the problem of explainable misinformation detection by picking up on minute contextual inconsistencies?
-We use detailed external information to inform the model’s predictions through the external retrieval module, which utilises reverse image-based search in order to provide the agents with external real-world context related to the image-text pair.
-We carry out our experiments with the GPT-4o (OpenAI ) model to achieve state-of-the-art performance on the misinformation detection task while also providing detailed and coherent explanations for the predictions.
-We achieve this without any domain-specific fine-tuning, thus ensuring easier and faster generalization to other domains in addition to low computational overhead.
+We present LLM-Consensus, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1).
+Unlike prior work, which largely provides predictions without explanations, our approach uses multiple multi-modal models debating to determine whether an image-text pair constitutes misinformation.
+We address the question:
+Can debating multi-modal models, equipped with external context, detect misinformation by identifying subtle contextual inconsistencies?
+Our external retrieval module uses reverse image search to provide agents with real-world context, enhancing their predictions.
+Using the GPT-4o (OpenAI, ) model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning.
+This ensures faster generalization to new domains with minimal computational overhead.
 3.1 Debate Modelling
 Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways.
-We explore multiple debating strategies to structure the conversation between agents.
+We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments.
 Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s).
@@
 Asynchronous Debate (not) against Human:
-We define an asynchronous debating strategy in which models wait for the other participants’ responses before generating their own.
-Figure 2 (a) and (b) show synchronous and asynchronous debating structures, respectively.
-While synchronous debates, where all participants speak at once, can be faster and computationally more efficient, we opt for an asynchronous setting where each model response is based on previous responses of other models.
-We observe that this method of structuring the debate works better since models are able to pick up on contextual ambiguities in their responses in a more organised and structured way which is crucial to the process of misinformation detection.
-An important point to note for this setup is the way we structure model prompts.
-The debating models are not aware that they are debating other AI agents.
-The prompts are structured such that each debating model believes it is talking to a human.
+One of the core setups we test in our experiments is an asynchronous debating strategy, where models wait for the responses of other participants before generating their own.
+Figure 2 (a) and (b) illustrate the differences between synchronous and asynchronous debate structures.
+While synchronous debates, where all participants respond simultaneously, can be faster and more computationally efficient, we find the asynchronous setup more effective.
+This structure allows models to better identify contextual ambiguities in a more organized and systematic manner, which is crucial for misinformation detection.
+A key aspect of this setup is the design of model prompts: the debating models are not aware that they are interacting with other AI agents but are instead led to believe they are debating with a human.
 Judged Debate:
@@
 Models are incentivised to structure their arguments in a way that makes them most convincing to the judge.
-We structure this debate configuration similar to Khan et al.
-(2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
+We structure this debate configuration similar to (Khan et al., 2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
 Actor-Skeptic:
@@
 Debate with Disambiguation:
-Improving on the actor-skeptic method, in this setup, we allow all agents to act as actors and skeptics.
-Models are tasked with not only generating their own responses but also disambiguation queries to refine further or refute the other agents’ responses.
-These disambiguation queries are then used to search the Internet to obtain information to refine model outputs further.
-We are the first to propose this debate setup and while it does not achieve the best results in this work, we believe future research can greatly benefit by refining this setup further.
-We empirically identify the best debate set up to be the asynchronous debate where the model believes it is debating against a human and not an AI agent.
+Building on the actor-skeptic method, we allow all agents to act as both actors and skeptics.
+Models generate their own responses and disambiguation queries to refine or challenge other agents’ outputs.
+These queries are used to retrieve additional information from the Internet, further improving model responses.
 3.2 Prompt Engineering
-The debate structure is substantiated through prompt engineering.
-Figure 1 shows that the first stage of our method requires for each AI agent to generate an independent response to whether the given image-text pair is misinformation.
-Each agent must take into account the external context related to the image obtained through the external information retrieval module.
-Specifications of the various prompts used in this work can be found in Appendix A.3.
-An initial prompt provides the agent with a summary of the news articles related to the image and, based on it, asks the agent to classify the image-text pair as misinformation or not.
-The prompt asks the agent to focus on certain details in the image, such as watermarks, flags, etc.
-We observe that images used in news articles often contain minute yet crucial details which can be used to inform the final decision about whether the image actually belongs to the news articles.
-Therefore, prompting the agent to pay special attention to these details further helps detect inconsistencies.
-Once the agents have formed independent opinions about the image-text pair, they must then participate in a debate.
-While the same prompt can be used for each debate round, we provide a different prompt for the first round to allow each agent to understand the changing nature of the conversation.
-Responses from other agents are provided as a part of the prompt, and the agent is asked to agree or disagree.
-The agent must also clearly state the reasoning process behind its argument.
-The prompt further requires the agent to identify ambiguities in the other agents’ reasoning.
-This allows the agent to closely analyse the responses from other agents and use them to inform its own decision.
-A separate prompt is then used to facilitate all rounds of the debate following round one.
-It takes into account the other agent’s suggestions from the previous round and, after refining its own response, asks the agent to point out any inconsistencies in this new response.
-We note that in such a scenario, agents are prone to simply agreeing with each other and repeating the other agent’s response.
-This is especially the case when agents believe they are conversing with a human.
-However, this tendency of the agents to easily give up their own opinions and simply agree with the other participants does not facilitate an advantageous debate, and the agents do not discover any new information related to the image-text pair.
-Therefore, the debate prompt also contains clear and explicit instructions asking the agent not to simply agree with the presented response unless it has an acceptable reason to do so.
-We find this helps the agents develop stronger stances and reluctance to blind agreement.
+The debate structure is enabled through prompt engineering.
+As shown in Figure 1, the first stage involves each AI agent independently determining whether the image-text pair is misinformation, using context from the external retrieval module (details in Appendix A.3).
+The initial prompt summarizes related news articles and instructs the agent to focus on key image details, like watermarks and flags, to identify inconsistencies.
+After forming initial opinions, the agents engage in a debate.
+The first round uses a unique prompt to help agents adapt to the discussion, while subsequent rounds use a standard prompt incorporating suggestions from previous rounds.
+Agents must agree or disagree with others’ responses while identifying ambiguities and refining their reasoning.
+To prevent blind agreement, the prompt explicitly instructs agents to provide valid reasoning before accepting others’ responses.
+This promotes stronger, independent stances and enhances the discovery of new information.
 3.3 External Information Retrieval
-Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al.
-(2023)).
+Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis, ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al., 2023).
 This summary is then passed to the debating agents as a part of the initial prompt.
+A model’s world knowledge is limited by its training data and time frame, but external retrieval allows access to up-to-date information beyond these constraints.
+Previous work relies on external datasets (Abdelnabi et al., 2022), but these often provide only news article titles, which lack sufficient detail.
 Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame).
-Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al.
-(2022)) to supplement external information related to an image-caption pair.
+Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al., 2022) to supplement external information related to an image-caption pair.
 However, we find this information lacking in detail since it is limited to the title of a news article.
-Agents can greatly benefit from the knowledge of the entire news article and its content rather than just the title when making a decision about whether a given image-caption pair, when considered in the context of the news article, is misinformation.
-To this end, we propose our own external information retrieval module.
-We observe a significant improvement in accuracy after incorporation of the external information retrieval module into the pipeline.
-The module is implemented in two stages:
+Full access to the article content can significantly enhance an agent’s ability to assess whether an image-caption pair constitutes misinformation.
+To address this, we propose a dedicated external information retrieval module, which improves accuracy when integrated into the pipeline.
+The module operates in two stages:
 3.3.1 API-Based Information Retrieval
-The Bing Visual Search API (vis ) is used for the task of obtaining web pages related to a given image.
+The Bing Visual Search API is used for the task of obtaining web pages related to a given image (vis, ) .
 A given image from the dataset is used to obtain a list of web pages completely and partially related to the image.
@@
 We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used.
-Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al.
-(2021)), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results.
+Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al., 2021), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results.
 In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base.
-Since this is not the case for a significant fraction of examples in the dataset, it does not adversely affect system performance.
+Since this is relevant for only a very small number of examples in the dataset, it does not adversely affect system performance.
 3.3.2 Summarization using LLM
-Once the top three web pages have been identified, we scrape the text from the web pages to obtain the textual information related to the context in which the image appears on the Internet.
-The compiled textual information is often too long to be passed directly to the agent, and hence, we use the Llama-13B (Touvron et al.
-(2023)) language model to summarize this information.
-The summaries obtained from the LLM only focus on the most important parts of the text and hence also allow agents to develop a more focused understanding of the external context.
-While this method of summarization works for most samples in our dataset, there are some examples where the obtained web pages are not in the English language, and the LLM struggles with summarization.
-In this regard, we add an additional check that ignores text from web pages in languages other than English.
-While this restricts our system to the English language, it does not adversely affect system performance due to the distribution of the dataset, which consists of images mostly taken from English-language news articles.
-Multi-lingual support can be achieved by first translating text to English and then summarizing it.
+Once the top three web pages are identified, we scrape their text to gather contextual information about the image.
+This textual data is often too lengthy for direct input to the agent, so we use the Llama-13B (Touvron et al., 2023) model to generate concise summaries, focusing on key details to help the agent better understand the external context.
+For non-English web pages, the LLM struggles with summarization, so we add a filter to exclude them.
+Although this restricts the system to English, it does not impact performance due to the dataset’s predominantly English-language content.
+Multi-lingual support could be added by translating text to English before summarization.
 3.4 Coherent Reasoning
-All the different components of MAD-Sherlock are brought together in this stage of the pipeline.
+All the different components of LLM-Consensus are brought together in this stage of the pipeline.
 Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same.
@@
 4.1 Dataset
-We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al.
-(2021)).
-The dataset is built based on the VisualNews (Liu et al.
-(2020)) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian and The Washington Post.
+We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al., 2021).
+The dataset is built based on the VisualNews (Liu et al., 2020) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian, and The Washington Post.
 The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair.
-CLIP (Radford et al.
-(2021)) is used to retrieve semantically similar images for a given caption.
+CLIP (Radford et al., 2021) is used to retrieve semantically similar images for a given caption.
 We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples.
 The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively.
+We use the NewsCLIPpings dataset for evaluating our model’s efficacy in identifying out-of-context images and information due to its well-established use in prior work, its relative scale compared to other available datasets, and the lack of more recent large-scale benchmarks in this domain.
+Though no single dataset can capture the full diversity of real-world misinformation, NewsCLIPpings provides a standardized evaluation setting that facilitates comparability with existing methods.
 4.2 Experimental Setup
-All experiments were run on 8 A40 (46GB) Nvidia GPU server.
+All experiments were run on 8 A40 (46GB) Nvidia GPU server. The estimated cost of processing one data sample using LLM-Consensus is \$0.24. Inference times range from 5 to 15 seconds.
 Debate Setup:
-We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al.
-(2024b)).
+We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al., 2024b).
 The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset.
@@
 External Retrieval Module:
-We use the Bing Visual Search API (vis ) to run an image-based reverse search.
-Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new ).
-Finally, we use Llama-13B (Touvron et al.
-(2023)) to summarise the text obtained from the top $k=3$ web pages.
+We use the Bing Visual Search API (vis, ) to run an image-based reverse search.
+Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new, ).
+Finally, we use Llama-13B (Touvron et al., 2023) to summarise the text obtained from the top $k=3$ web pages.
 This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
 Baselines and Prior Work:
-We compare MAD-Sherlock to existing pretrained multi-modal baselines including CLIP (Radford et al.
-(2021)), VisualBERT (Li et al.
-(2019)), InstructBLIP (Dai et al.
-(2023)) and LLaVA (Liu et al.
-(2024b)).
-We also compare performance against GPT-4o (OpenAI & et al.
-(2024); OpenAI ).
+We compare LLM-Consensus to existing pretrained multi-modal baselines including CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2024b).
+We also compare performance against GPT-4o (OpenAI & et al., 2024; OpenAI, ).
 The models are presented with the image and caption pair and asked if the pair is misinformation.
 The models are further prompted to explain their reasoning.
-We also show results for two baseline methods trained from scratch, namely EANN (Wang et al.
-(2018)) and SAFE (Zhou et al.
-(2020)).
-We further compare MAD-Sherlock to DT-Transformer (Papadopoulos et al.
-(2023)), CCN (Abdelnabi et al.
-(2022)), Sniffer (Qi et al.
-(2024)), VINVL (Huang et al.
-(2024)), SSDL (Mu et al.
-(2023)) and Neuro-Sym (Zhu et al.
-(2022)).
+We also show results for two baseline methods trained from scratch, namely EANN (Wang et al., 2018) and SAFE (Zhou et al., 2020).
+We further compare LLM-Consensus to DT-Transformer (Papadopoulos et al., 2023), CCN (Abdelnabi et al., 2022), Sniffer (Qi et al., 2024), VINVL (Huang et al., 2024), SSDL (Mu et al., 2023) and Neuro-Sym (Zhu et al., 2022).
 4.3 Results
-We present results for the experiments conducted to select the best debate setup as well compare the performance of MAD-Sherlock against existing methods.
+We present results for the experiments conducted to select the best debate setup as well compare the performance of LLM-Consensus against existing methods.
 We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
@@
 Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations.
-We also report results for the Asynchronous Debate setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain.
+We also report results for the Async_Debatehuman setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain.
 The external retrieval of information significantly boosts performance.
-All following debate set-ups, therefore, use external information related to the image-caption pair as a part of their initial prompt.
 We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent.
 Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human.
-Further, the Asynchronous Debate setup benefits from the ensemble of agents which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses.
+Further, the asynchronous debate setup benefits from the ensemble of agents, which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses.
 The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought.
@@
 We present our results on the NewsCLIPpings dataset against existing out-of-context detection methods discussed in section 4.2.
-Model Accuracy
+Model Accuracy $\uparrow$
 SAFE 50.7
@@
 Neuro-Sym 68.2
+GPT-4o# (w retrieval) 86.0
 Sniffer (w finetuning) 88.4
 Sniffer (w/o finetuning) 84.5
-MAD-Sherlock (ours) 90.17
-Table 2: Performance comparison between our model and baselines: MAD-Sherlock (with GPT-4o) out performs all related work.
+LLM-Consensus (ours) 90.8
+Table 2: Performance comparison between our model and baselines: LLM-Consensus (with GPT-4o) out performs all related work.
+Note: the GPT-4o# setup is identical to LLM-Consensus with the absence of a multi-agent debate, here only a single agent with access to external information is considered.
 Table 2 shows the comparison between our system and existing methods.
-We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al.
-(2024); OpenAI ) model.
+We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al., 2024; OpenAI, ) model.
 Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset.
-While we do not provide a quantitative assessment of explanations by MAD-Sherlock, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines.
+While we do not provide a quantitative assessment of explanations by LLM-Consensus, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines.
+This is attributed to the fact that in a multi-agent setup, we have multiple context windows which leads to more coherent and relevant final explanations.
 We leave the detailed analysis of these explanations and the development of the associated metrics as future work.
+We also note that the debate paradigm in itself is essential to the system performance.
+We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model.
 We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work.
@@
 These models require additional integration into more comprehensive pipelines, as done in this work.
-In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al.
-(2020)) and EANN (Wang et al.
-(2018)) perform worse than pretrained multi-modal models.
+In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al., 2020) and EANN (Wang et al., 2018) perform worse than pretrained multi-modal models.
 This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
+4.3.3 Ablations
+To analyze the importance of each component of the LLM-Consensus framework, we conduct ablation experiments.
+Specifically, we evaluate the effect of using LLaVA against GPT-4o, the impact of the external retrieval module, and the power of the multi-agent debate framework.
+LLaVA GPT-4o Retrieval Debate Accuracy Precision Recall
+$77.1$ $68.4$ $89.3$
+$86.2$ $82.6$ $90.6$
+$86.0$ $80.2$ $95.6$
+$90.2$ $\mathbf{90.3}$ $90.1$
+$\mathbf{90.8}$ $85.5$ $\mathbf{99.0}$
+Table 3: Ablation: Quantitative evaluation of each component of LLM-Consensus on classification performance.
+We observe that the combination of GPT-4o, the external retrieval module, and the multi-agent debate framework yields the highest performance across all metrics, with 90.8% accuracy, 85.5% precision, and 99.0% recall, demonstrating the value of combining these components.
+The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7 (as seen in Table 2) to 90.2, underscoring its role in enabling contextual reasoning and refining predictions.
+Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3 to 85.5 and recall from 90.1 to 99.0.
+Meanwhile, retrieval contributes more substantially to LLaVA’s performance gains, likely due to GPT-4o’s broader world knowledge.
+Without external retrieval or the debate framework, the performance drops, emphasizing the critical role of these components in achieving state-of-the-art results.
 5 User Study
 We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation.
-While it is easy to quantify model performance in terms of misinformation detection, there are no effective metrics to assess the quality of the explanations generated by the model.
-Therefore, in order to perform a thorough analysis of the system performance, a user study is essential.
-For a deeper analysis we further grouped the participants based on their profession into three groups, namely: Journalists, AI Academics (studying AI) and Others.
-Further details regarding the study setup and participant groups can be found in Appendix A.4.
-In the study, participants were shown ten image-text pairs and were asked to decide if the image and caption, when considered together, were misinformation or not.
-They were also asked to provide a confidence rating for their answer on a scale of 0-10, with 10 being the highest confidence level.
-For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair.
-These AI insights were the final output explanations from MAD-Sherlock.
-Participants were then asked to reconsider their answers in light of the new information from the AI agent.
-Table 3 shows that average system performance is better than the average human performance for both cases where the participants have access to AI insights and where they do not.
-Therefore, MAD-Sherlock can be used as a reliable assistive tool for OSINT research for detecting and explaining misinformation with little or no human intervention.
-Study Setup Average Accuracy
+While model performance in misinformation detection is easy to quantify, there are no established metrics for assessing the quality of its explanations, making the user study essential.
+Participants were grouped into three categories based on their professions: Journalists, AI Academics, and Others (see Appendix A.4 for details).
+During the study, participants reviewed ten image-text pairs and decided whether each pair constituted misinformation.
+They also rated their confidence on a scale of 0 to 10.
+After submitting their initial answers, participants were shown LLM-Consensus’s AI-generated explanations and asked to reconsider their responses.
+As shown in Table 4, the system outperformed the average human performance in both cases, with and without AI insights.
+These results demonstrate LLM-Consensus’s potential as a reliable assistive tool for OSINT research, capable of detecting and explaining misinformation with minimal human intervention.
+[…]
 Humans $60.3\pm 13.5$
-Humans+MAD-Sherlock $76.7\pm 12.2$
-MAD-Sherlock $\mathbf{80.0\pm 0.0}$
-Table 3: Performance comparison between different study setups: MAD-Sherlock outperforms humans with and without AI assistance.
-We further observe that the average human accuracy for the misinformation detection task increases by more than 27% concretizing the fact that AI insights from our model do actually improve human efficiency in detecting misinformation.
-We also observe interesting patterns for group-wise analysis which we believe would be valuable for future work.
-Table 4 shows that the performance of all groups improves significantly and is not far off from that of professional journalists.
-The average confidence level (out of 10) is comparable across all the groups before and after considering MAD-Sherlock insights and generally increases.
-Therefore, we conclude that MAD-Sherlock can significantly uplift non-expert performance and hence can be useful in citizen intelligence applications.
-Avg acc $\uparrow$ Avg conf $\uparrow$ Avg acc $\uparrow$ Avg conf $\uparrow$
-Group (only human) (only human) (with (with
-MAD-Sherlock) MAD-Sherlock)
-Journalists $70.0\pm 1.4$ $4.3\pm 2.1$ $82.2\pm 0.9$ $5.3\pm 1.3$
-AI Academics $60.7\pm 1.4$ $3.2\pm 0.8$ $79.3\pm 1.3$ $5.8\pm 1.4$
-Others $56.7\pm 1.5$ $3.9\pm 1.2$ $71.7\pm 1.1$ $5.8\pm 1.4$
-Table 4: Performance comparison between different participant groups: All groups show performance improvement with MAD-Sherlock.
-AI Academics are able to perform nearly at par with professional journalists after considering insights from MAD-Sherlock.
+Humans+LLM-Consensus $76.7\pm 12.2$
+LLM-Consensus $\mathbf{80.0\pm 0.0}$
+Table 4: Performance comparison between different study setups: LLM-Consensus outperforms humans with and without AI assistance.
+Group-wise analysis, shown in Table 5, reveals significant performance gains across all groups, with results approaching those of professional journalists.
+Confidence levels (out of 10) are comparable across groups and generally increase after using LLM-Consensus insights.
+Thus, LLM-Consensus can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
+Metric Journalists Academics Others
+Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
+Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
+Accuracy (with LLM-Consensus) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
+Confidence (with LLM-Consensus) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
+Table 5: Performance comparison: LLM-Consensus improves performance across all participant groups.
 6 Conclusion and Future Work
@@
 Agents in such a setting show enhanced abilities to critically evaluate an argument and pick up on minute inconsistencies.
-Our final system, MAD-Sherlock, achieves state-of-the-art performance on the misinformation detection task.
-Further, owing to our advanced external retrieval module, MAD-Sherlock provides clear, coherent and detailed explanations.
-As a result, MAD-Sherlock significantly improves the OOC misinformation detection performance of both human experts, and non-experts.
-We identify several promising avenues for future research in this field.
-Notably, improving how agents handle disambiguation queries could significantly enhance detection accuracy and explainability by providing access to additional information.
-The research community would benefit from a continuously updated benchmark dataset, incorporating more recent news articles and subtler inconsistencies.
-A direct extension of this work involves applying our methods to video-text pairs.
-Since MAD-Sherlock’s explanations are currently limited to text, future research should explore multi-modal and mixed-modal explanation generation.
-Additionally, further experiments with different hyper-parameters, such as varying the number of agents and debate rounds, could refine the effectiveness of our approach.
-Finally, while we conducted extensive user studies with MAD-Sherlock, deploying it on a larger scale in professional environments and within the citizen intelligence community will provide valuable insights into its real-world performance, uncovering new opportunities for improvement.
-For an analysis of limitations, please refer to Appendix A.1.
-7 Acknowledgments
-A part of this work was supported by UKRI DECaDE EP/T022485/1 and the UKRI grant: Turing AI Fellowship EP/W002981/1, Armasuisse Science+Technology.
-We would also like to thank the Royal Academy of Engineering and FiveAI.
+Our final system, LLM-Consensus, achieves state-of-the-art performance in misinformation detection while offering clear, detailed explanations, thanks to our advanced external retrieval module.
+As a result, we observe significant improvements in out-of-context (OOC) detection for both experts and non-experts.
+We identify several promising directions for future research.
+The community would benefit from a continuously updated benchmark dataset incorporating recent news articles and subtler inconsistencies.
+We plan to extend our methods to video-text pairs and multi-lingual content.
+We also believe that integrating more advanced models in the summarization pipeline will further enhance performance and are interested in comparing LLM-Consensus with multi-agent systems using external retrieval.
+Finally, we conducted extensive user studies but recognize the importance of large-scale deployment in professional settings and the citizen intelligence community.
+This will provide valuable real-world feedback and uncover further opportunities for improvement.
+For a discussion of limitations, see Appendix A.1.
 Appendix A Appendix
 A.1 Limitations
-Despite the strong performance of MAD-Sherlock, several limitations remain.
+Despite the strong performance of LLM-Consensus, several limitations remain.
 First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve.
@@
 Finally, our dataset, though comprehensive, primarily focuses on English-language news, limiting the generalizability of the system across non-English contexts.
-Another important limitation is the potential risk that open-sourcing MAD-Sherlock might allow adversaries to train models specifically designed to counter or evade detection by our system.
+Another important limitation is the potential risk that open-sourcing LLM-Consensus might allow adversaries to train models specifically designed to counter or evade detection by our system.
 As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness.
@@
 Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
-Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining MAD-Sherlock’s effectiveness in the long term.
+Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining LLM-Consensus’s effectiveness in the long term.
 The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
-Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of MAD-Sherlock.
+Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of LLM-Consensus.
 A.2 Sample Image-Caption Pair in the News Domain
@@
 Image and caption taken from the BBC article here (Accessed at 17:43 on Aug 11, 2024): https://www.bbc.co.uk/news/articles/cze5pkg5jwlo
-A.3 Prompts for MAD-Sherlock
+A.3 Prompts for LLM-Consensus
 This is a summary of news articles related to the image: {}
@@
 For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair.
-These AI insights were the final outputs from MAD-Sherlock.
+These AI insights were the final outputs from LLM-Consensus.
 Participants were then asked to reconsider their answer and again decide if the image-text pair was misinformation or not, in light of the new information from the AI agent.
@@
 This was done to ensure an unbiased estimate of average human performance.
-The image-text pairs to include in the study were taken from the NewsCLIPpings (Luo et al.
-(2021)) dataset.
+The image-text pairs to include in the study were taken from the NewsCLIPpings (Luo et al., 2021) dataset.
 AI insights were taken from our best-performing setup involving the GPT-4o model.
@@
 Therefore the model accuracy for the task was 80% and we use this as the baseline accuracy to compare human performance against.
-We analyse two special cases, where MAD-Sherlock argues for the wrong answer.
+We analyse two special cases, where LLM-Consensus argues for the wrong answer.
 We include these results in order to observe how persuasive our system can be even when it is wrong.
-We note in the instance where the image-text pair was actually misinformation and the model argued that it was not, 6 participants changed their correct responses to those suggested by MAD-Sherlock.
+We note in the instance where the image-text pair was actually misinformation and the model argued that it was not, 6 participants changed their correct responses to those suggested by LLM-Consensus.
 Although this is only 5% of the participants, it still gives a significant insight into how persuasive the model can appear even when it is wrong.
 While the case of false negatives is important, false positives are an even more concerning matter for our problem statement.
-In the case where MAD-Sherlock declared the given image-text pair to be misinformation when it was not, is important to analyse.
+In the case where LLM-Consensus declared the given image-text pair to be misinformation when it was not, is important to analyse.
 In this setting 50% of the total participants changed their answer to the wrong one, therefore believing a piece of true information to be false.
@@
 The same task without AI insights would require extensive analysis and we project it would take between 30-45 minutes to decide if 10 image-text pairs were misinformation.
-Acknowledgements
-Part of this work was supported by UKRI DECaDE EP/T022485/1
+A.5 Multi-modal debates for hamful meme detection
+While this work relates to a different problem than OOC misinformation detection in the news domain, we still find the approach taken by the authors a relevant related work and therefore include it here.
+(Lin et al., 2024) use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful.
+These explanations are then used to train a small language model as a judge to determine whether the image and text that make up the meme are actually harmful.
+This work does not allow agents to have flexibility of opinion.
+There are always two agents, and each one is provided a stance to defend.
+Moreover, a judge decides the final outcome of the debate and needs to be trained on data from the debate.
+This method also does not benefit from external retrieval, and therefore, the debating agents are not aware of the crucial external context related to the input.
+Finally, this work is related to harmful meme detection and does not concern the problem of misinformation detection in the news domain, which likely requires more intricate contextual analysis, including of external context.

### v3

@@
-LLM-Consensus: Multi-Agent Debate for Visual Misinformation Detection
+MAD-Sherlock: Multi-Agent Debate for Visual Misinformation Detection
 Abstract
-One of the most challenging forms of misinformation involves the out-of-context (OOC) use of images paired with misleading text, creating false narratives.
-Existing AI-driven detection systems lack explainability and require expensive finetuning.
-We address these issues with LLM-Consensus, a multi-agent debate system for OOC misinformation detection.
-LLM-Consensus introduces a novel multi-agent debate framework where multimodal agents collaborate to assess contextual consistency and request external information to enhance cross-context reasoning and decision-making.
-Our framework enables explainable detection with state-of-the-art accuracy even without domain-specific fine-tuning.
-Extensive ablation studies confirm that external retrieval significantly improves detection accuracy, and user studies demonstrate that LLM-Consensus boosts performance for both experts and non-experts.
-These results position LLM-Consensus as a powerful tool for autonomous and citizen intelligence applications.
-misinformation detection,out-of-context image use,LLMs,multimodal models,multi-agent debate,safety
+One of the most challenging forms of misinformation involves pairing images with misleading text to create false narratives.
+Existing AI-driven detection systems often require domain-specific finetuning, limiting generalizability, and offer little insight into their decisions, hindering trust and adoption.
+We introduce MAD-Sherlock, a multi-agent debate system for out-of-context misinformation detection.
+MAD-Sherlock frames detection as a multi-agent debate, reflecting the diverse and conflicting discourse found online.
+Multimodal agents collaborate to assess contextual consistency and retrieve external information to support cross-context reasoning.
+Our framework is domain- and time-agnostic—requiring no finetuning–yet achieves state-of-the-art accuracy with in-depth explanations.
+Evaluated on NewsCLIPpings, VERITE, and MMFakeBench, it outperforms prior methods by 2%, 3%, and 5%, respectively.
+Ablation and user studies show that the debate and resultant explanations significantly improve detection performance and improves trust for both experts and non-experts, positioning MAD-Sherlock as a robust tool for autonomous citizen intelligence.
+multi-agent debate,agents,misinformation
 1 Introduction
-Our growing dependence on online channels for news and social networking has been complemented by a surge in exploits of digital misinformation (Aslett et al., 2024; Hasher et al., 1977; Brashier & Marsh, 2020).
-While many manipulation techniques pose serious threats, one of the most prevalent methods for creating fake online content is the out-of-context (OOC) use of images (pbs, ).
-This involves using unaltered images in a misleading, false context to convey deceptive information, a strategy that requires minimal technical expertise.
-OOC misinformation detection requires a nuanced understanding of the relationship between text and images and the ability to identify when they are misaligned.
-Detecting these subtle inconsistencies is time-consuming for humans, and (Sultan et al., 2022) shows that time pressure further reduces detection accuracy, exacerbating scalability issues.
-As a result, AI-driven tools have gained attention for scaling detection efforts.
-However, conventional deep learning forensic techniques (Castillo Camacho & Wang, 2021; Heidari et al., 2024; Zhu et al., 2018; Amerini et al., 2021; Hina et al., 2021), designed to detect manipulations like PhotoShop editing (Tolosana et al., 2020; Masood et al., 2023; Farid, 2016; Wang et al., 2019) and AI-generated Deepfakes (mit, ), focus on spotting artifacts from tampering.
-In contrast, OOC detection requires cross-contextual reasoning, as the deception stems from the misalignment of legitimate images with misleading text.
-Figure 1: Overview of LLM-Consensus: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not.
+The rise of online news and social media has been paralleled by a surge in digital misinformation [4, 14, 6].
+Among the most widespread tactics is out-of-context (OOC) image use [13], where unaltered images are paired with misleading text to deceive, requiring little technical skill.
+OOC detection demands nuanced reasoning to identify misalignment between images and accompanying text.
+This task is time-consuming for humans, and detection accuracy drops under time pressure [37], limiting scalability.
+AI tools offer a path forward, but traditional forensic methods [8, 15, 46, 2, 16] focus on tampering artifacts (e.g., Photoshop edits [12, 40] or Deepfakes [38, 26, 36]) and are ill-suited for OOC detection, which hinges on cross-contextual reasoning rather than visual manipulation.
+Figure 1: Overview of MAD-Sherlock: Two or more independent agents see the same image-text input and are tasked with detecting whether the input is misinformation or not.
 After the agents form their independent opinions, they participate in a debate until they converge on the same response or when $n$ debate rounds are completed (whichever is earlier).
-Pretrained Large Multimodal Models  (Liu et al., 2024b; OpenAI & et al., 2024; Li et al., 2019; Radford et al., 2021, LMMs) provide a promising direction for detecting OOC use of images for their ability to process both text and image content in tandem.
-However, using LMMs directly for OOC detection presents several challenges, particularly in the news domain.
-For instance, news articles often include images that are not directly related to the article’s content.
-An article about the 2024 U.S.
-presidential candidates, for example, might feature a close-up of Donald Trump from an unrelated online database.
-Although the image was taken outside the election period, it is not considered OOC since it doesn’t misrepresent the article’s context.
-Such cases complicate LMMs’ ability to accurately identify OOC usage based solely on their pre-trained knowledge, as this knowledge may be outdated or insufficiently detailed.
-Moreover, even with recent advancements, LMMs are capable of hallucinating and, hence, generating false information (Bai et al., 2024; Liu et al., 2024a).
-While rapidly improving, they sometimes fail to understand user instructions and intent correctly.
-We show that off-the-shelf LMMs indeed suffer from these issues, reducing their ability to detect OOC misinformation in practice.
-While prior work (Qi et al., 2024) has shown that off-the-shelf models can be improved using task-specific fine-tuning, this approach is resource-intensive and requires continual updating to keep up with recent events.
-Moreover, detecting OOC images only solves part of the problem.
-The real value lies in being able to explain the OOC use of pictures in human-readable form.
-It can be instrumental for human validators to observe the model’s line of logic and gain better insight into, and trust in, the classification process.
-In this work, we propose a novel LMM-based post-training approach for scalable OOC misinformation detection that simultaneously improves contextual reasoning, provides in-built explainability, and achieves state-of-the-art detection accuracy even without task-specific fine-tuning (see Section 3).
-Specifically, our framework LLM-Consensus frames the detection problem as a dialectic debate between multiple LMM agents, where, in contrast to prior work (Minsky, 1988; Li et al., 2023a; Du et al., 2023a; Khan et al., 2024)), agents have access to external information retrieval.
-Compared to single-agent chain-of-thought approaches (Wei et al., 2024), the use of multiple agents allows for a clean separation of agent contexts, decentralisation of action spaces, and opportunities for parallel computation (Schroeder de Witt et al., 2020; Du et al., 2023b).
-In addition, due to its compositional nature, both additional human and autonomous agents can be dynamically added to the multi-agent reasoning process, allowing the use of LLM-Consensus as an interactive tool for human experts.
-To the best of our knowledge, no prior work has used debating LMMs for detecting and explaining OOC image use.
-We perform a comprehensive empirical evaluation of our method (see Section 4), including the study of multiple debate configurations.
-To optimise OpenAI API use, we utilize an experimental pipeline where preliminary experiments are performed using the open-source LLaVA model (Liu et al., 2024b), which is only later replaced with GPT-4o (OpenAI, ) to achieve state-of-the-art performance.
-We find that LLM-Consensus outperforms both prior work and novel baselines that we introduce, is more robust to various failure modes, and produces coherent explanations that help both human experts and non-experts significantly improve their detection accuracy in user studies.
-We identify both access to external information retrieval and complete freedom of opinion as key ingredients to the performance of LLM-Consensus.
-Finally, we discuss current limitations of our method and propose future work toward overcoming the scalability challenges in large-scale online OOC misinformation detection.
+Pretrained large multimodal models (LMMs) offer a promising foundation for detecting OOC image use by jointly interpreting text and images [23, 30, 19, 34].
+However, applying them to news content is challenging.
+News articles often feature loosely related but contextually appropriate images—e.g., a pre-2024 photo of Donald Trump used in election coverage—which makes OOC detection difficult using pretraining alone.
+LMMs also suffer from hallucinations, misinterpret intent [5, 22], and lack up-to-date knowledge.
+Off-the-shelf models exhibit these limitations, reducing their reliability.
+Fine-tuning helps [33], but is costly and requires frequent updates.
+Crucially, beyond detection, models must provide clear, human-readable explanations to support understanding and trust in their decisions.
+In this work, we propose a novel post-training framework for scalable OOC misinformation detection that improves contextual reasoning, offers built-in explainability, and achieves state-of-the-art accuracy without task-specific fine-tuning (see Section 3).
+Our method frames the detection problem as a dialectic debate between LMM agents, augmented with external information retrieval.
+Unlike single-agent chain-of-thought methods [42], our multi-agent setup enables context separation, decentralized reasoning, and parallelization [35, 11].
+Prior work [27] shows agents struggle to maintain diverse perspectives within a single context window; our approach addresses this via modular, compositional interaction.
+It also allows seamless integration of human or autonomous agents, making MAD-Sherlock a flexible tool for expert oversight.
+To our knowledge, this is the first use of multi-agent LMM debate for both detecting and explaining OOC image use.
+MAD-Sherlock is backbone-agnostic, compatible with open- and closed-source models.
+We prototype and tune with LLaVA [23] to reduce API costs, then deploy GPT-4o [29] for final results.
+By avoiding task-specific fine-tuning, our method remains broadly applicable across domains and time periods.
+We evaluate on three benchmarks—NewsCLIPpings, VERITE, and MMFakeBench—and achieve new state-of-the-art performance across all.
+MAD-Sherlock outperforms prior methods and baselines, demonstrates robustness, and produces detailed, interpretable explanations.
+In user studies, these explanations significantly improve detection accuracy for both experts and non-experts.
+Key contributors to performance include external retrieval and agent independence. We discuss limitations and outline future directions toward scalable, general-purpose AI for public good.
 2 Related Work
-Recent work has focused on using joint image-text representations to classify an instance as OOC.
-(Aneja et al., 2022) follow a self-supervised approach to assess whether two captions accompanying an image are contextually similar.
-They enforce image-text matching during training by formulating a scoring function to align objects in the image with the caption.
-During inference, they use the semantic similarity between the two captions to classify them as OOC or not.
-The increased reliance on textual content limits the capabilities of this approach.
-This work also does not provide explanations for model predictions and is, therefore not interpretable.
-Moreover, this method works for image caption pairs where captions have information about objects in the image.
-This is not always the case with news articles (our domain of application), where captions can often just be related to the main content of an article rather than precisely describing the objects in the image.
-Appendix  A.2 shows an example of the same.
-(Abdelnabi et al., 2022) introduce the Consistency Checking Network (CCN), which models aspects of human reasoning across modalities for misinformation detection using external evidence aggregated from the Internet.
-The CCN employs memory networks to evaluate the consistency of image-caption pairs against retrieved evidence and a CLIP component (Radford et al., 2021) to assess the alignment between the image and caption.
-The use of external evidence enhances CCN’s classification performance compared to other methods.
-However, it lacks an explainability component, providing only binary decisions without further explanation.
-(Zhang et al., 2024) extend the neural symbolic method (Yi et al., 2019; Zhu et al., 2022) to develop an interpretable cross-modal misinformation detection model that provides supporting evidence for its predictions.
-They use symbolic graphs based on Abstract Meaning Representation (Banarescu et al., 2013) to detect out-of-context image use.
-Similarly, (Zhou et al., 2020) introduce Similarity Aware Fake news detection (SAFE), where neural networks jointly learn and analyze features and relationships between text and visual news representations to predict fake news.
-(Wang et al., 2018) introduce EANN: Event Adversarial Neural Networks to derive event invariant features which can be used to detect fake news that has recently been generated.
-EANN uses adversarial training to learn multi-modal features independent of news events.
-These methods require pretraining from scratch and, therefore, don’t benefit from the advanced reasoning capabilities and world knowledge of large pretrained models.
-Shalabi et al.
-(2023) use synthetic multi-modal data to establish the authenticity of image-text pairs.
-They use BLIP-2 (Li et al., 2023b) to generate a caption for the original image and Stable Diffusion (Rombach et al., 2022) to generate an image for the given original caption.
-This synthetic data is then used to reason that if the original image and caption are OOC, then the original and generated images should also be OOC as well as the original and generated text.
-This method relies on synthetic multi-modal data generation, which not only adds an additional computational overhead but also increases dependence on often unreliable synthetically generated data.
-Therefore, this method can suffer from issues related to generation models, including potential biases that these models may possess.
-This method also lacks interpretability.
-Sniffer (Qi et al., 2024) is the closest to our work.
-It uses the InstructBLIP (Dai et al., 2023) model to detect OOC image use and provide an explanation for its prediction.
-It makes use of internal and external knowledge using entity extraction APIs and image-based web searches.
-Information from all the sources is given to an LLM to predict and explain if an image has been used OOC.
-Sniffer only uses basic textual information, such as news article titles from websites, to form its external knowledge base.
-It also requires extensive training to adapt the model to the news domain which adds additional computational overhead and also restricts the generalization abilities of the model to other domains.
+Recent work has explored joint image-text representations for OOC classification.
+Aneja et al.
+[3] use a self-supervised approach to enforce image-text alignment via an object-caption matching score, classifying OOC instances by caption similarity.
+However, the method is heavily text-dependent and lacks interpretability.
+Abdelnabi et al.
+[1] introduce the Consistency Checking Network (CCN), which leverages memory networks and CLIP for image-caption consistency against external evidence, improving accuracy but offering only binary outputs without explanations.
+Zhang et al.
+[43] employ AMR-based symbolic graphs for interpretable OOC detection with evidence.
+Similarly, Zhou et al.
+[44] propose SAFE for joint text-visual reasoning in fake news detection.
+Wang et al.
+[41] present EANN, which uses adversarial training to extract event-invariant multimodal features.
+These methods require training from scratch and lack the advanced reasoning and knowledge of large pretrained models.
+Sniffer [33] is most similar to our work.
+It uses InstructBLIP [9] to detect OOC image use and generate explanations by aggregating internal and external knowledge from entity extraction APIs and image-based web searches.
+However, its external knowledge is limited to basic textual content (e.g., article titles), and adapting it to the news domain requires extensive training, reducing generalizability and increasing computational overhead.
+The CRAVE framework [10] clusters retrieved multimodal evidence into coherent narratives, then uses an LLM to produce fact-checking judgments with natural language explanations.
+While effective, treating all clusters equally may amplify fringe perspectives and distort the evidence base.
+Liu et al.
+[24] introduce MMD-Agent, a single-agent framework that hierarchically decomposes misinformation detection into textual, visual, and cross-modal subtasks before reasoning over their outputs to make a final prediction.
 3 Methodology
-We present LLM-Consensus, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1).
+We present MAD-Sherlock, an explainable misinformation detection system that jointly predicts and explains instances of misinformation (Figure 1).
 Unlike prior work, which largely provides predictions without explanations, our approach uses multiple multi-modal models debating to determine whether an image-text pair constitutes misinformation.
@@
 Our external retrieval module uses reverse image search to provide agents with real-world context, enhancing their predictions.
-Using the GPT-4o (OpenAI, ) model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning.
+Using the GPT-4o [29] model, we achieve state-of-the-art performance with detailed, coherent explanations, without requiring domain-specific fine-tuning.
 This ensures faster generalization to new domains with minimal computational overhead.
@@
 Analogous to real-world conversations, communication between two AI agents can also be structured in a myriad of ways.
-We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments.
-Instead of simple back-and-forth conversations, we opt for a debating set-up in which agents are asked to frame their own opinions and then defend them to other agent(s).
-We observe this facilitates more involved and detailed discussions among the models.
-Figure 2: Debating Strategies: We experiment with multiple debating strategies.
-The asynchronous debate setup where agents argue one after the other and take turns presenting their arguments is the best configuration.
+We explore multiple debating strategies to structure the conversation between agents, all of which are tested and evaluated in our experiments and informed by the work of Brown-Cohen et al.
+[7].
+Figure 2: Debating Strategies: We experiment with multiple debating strategies to evaluate which performs best on our task.
 Asynchronous Debate (not) Against Human:
-One of the core setups we test in our experiments is an asynchronous debating strategy, where models wait for the responses of other participants before generating their own.
-Figure 2 (a) and (b) illustrate the differences between synchronous and asynchronous debate structures.
-While synchronous debates, where all participants respond simultaneously, can be faster and more computationally efficient, we find the asynchronous setup more effective.
-This structure allows models to better identify contextual ambiguities in a more organized and systematic manner, which is crucial for misinformation detection.
-A key aspect of this setup is the design of model prompts: the debating models are not aware that they are interacting with other AI agents but are instead led to believe they are debating with a human.
+We experiment with an asynchronous debating strategy, where models wait for others’ responses before generating their own (Figure 2a vs.
+b).
+While synchronous debates are faster and more efficient, the asynchronous setup proves more effective for identifying contextual ambiguities—critical in misinformation detection.
+Notably, models are prompted to believe they are debating a human rather than another AI.
 Judged Debate:
-We also experiment with an asynchronous debate setup with a judge.
-Figure 2 (c) shows this setup.
-In this setup, models participate in an asynchronous debate as usual however, the final decision is made by a judge at the end of the debate.
-Models are incentivised to structure their arguments in a way that makes them most convincing to the judge.
-We structure this debate configuration similar to (Khan et al., 2024), where the judge does not have access to all the external information and has to rely only on the debate transcript to decide the final answer.
+We also explore an asynchronous debate setup with a judge (Figure 2c), where models debate as usual, but a final decision is made by a judge based solely on the debate transcript.
+Following Khan et al.
+[18], the judge lacks access to external information, incentivizing models to structure arguments that are maximally persuasive.
 Actor-Skeptic:
-In this setup, only one agent, the actor, is tasked with deciding whether a given image-text pair is misinformation.
-The agent generates a response which a skeptic then evaluates.
-The skeptic is tasked with finding logical errors in the actor’s argument and asking follow-up questions to disambiguate the actor’s response.
-It is important to note that neither the skeptic nor the actor has access to the ground truth.
-This setup does not benefit from an ensemble since both models assume different roles and only one agent is tasked with generating the final answer.
-Debate with Disambiguation:
-Building on the actor-skeptic method, we allow all agents to act as both actors and skeptics.
-Models generate their own responses and disambiguation queries to refine or challenge other agents’ outputs.
-These queries are used to retrieve additional information from the Internet, further improving model responses.
+In this setup, a single actor determines whether an image-text pair constitutes misinformation.
+A skeptic then critiques the actor’s reasoning, probing for logical flaws and ambiguities.
+Neither agent has access to the ground truth, and since their roles are distinct, this configuration does not benefit from ensembling.
 3.2 Prompt Engineering
-The debate structure is enabled through prompt engineering.
-As shown in Figure 1, the first stage involves each AI agent independently determining whether the image-text pair is misinformation, using context from the external retrieval module (details in Appendix A.3).
-The initial prompt summarizes related news articles and instructs the agent to focus on key image details, like watermarks and flags, to identify inconsistencies.
-After forming initial opinions, the agents engage in a debate.
-The first round uses a unique prompt to help agents adapt to the discussion, while subsequent rounds use a standard prompt incorporating suggestions from previous rounds.
-Agents must agree or disagree with others’ responses while identifying ambiguities and refining their reasoning.
-To prevent blind agreement, the prompt explicitly instructs agents to provide valid reasoning before accepting others’ responses.
-This promotes stronger, independent stances and enhances the discovery of new information.
+The debate is structured via prompt engineering (Figure 1).
+In the first stage, each agent independently assesses whether the image-text pair is misinformation, using context from an external retrieval module (see Appendix A.3).
+Prompts summarize relevant articles and emphasize visual cues (e.g., watermarks, flags) to guide initial judgments.
+Agents then debate: the first round uses a tailored prompt to initiate discussion; later rounds use a standard prompt incorporating prior responses.
+Agents must agree or disagree with peers while refining reasoning, and prompts explicitly discourage blind agreement by requiring justification for any alignment.
 3.3 External Information Retrieval
-Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API (vis, ) to obtain web pages related to a given image, which are then summarised using Llama-13B (Touvron et al., 2023).
+Figure 3: Structure of the external information retrieval module: We use the Bing Visual Search API to obtain web pages related to a given image, which are then summarised using Llama-13B [39].
 This summary is then passed to the debating agents as a part of the initial prompt.
-A model’s world knowledge is limited by its training data and time frame, but external retrieval allows access to up-to-date information beyond these constraints.
-Previous work relies on external datasets (Abdelnabi et al., 2022), but these often provide only news article titles, which lack sufficient detail.
-Since a model’s world knowledge is limited to its training data (and hence a particular time frame), incorporating external retrieval allows the model to access information beyond this training data (and time frame).
-Previous work makes use of pre-existing external retrieval-based datasets (Abdelnabi et al., 2022) to supplement external information related to an image-caption pair.
-However, we find this information lacking in detail since it is limited to the title of a news article.
-Full access to the article content can significantly enhance an agent’s ability to assess whether an image-caption pair constitutes misinformation.
-To address this, we propose a dedicated external information retrieval module, which improves accuracy when integrated into the pipeline.
-The module operates in two stages:
-3.3.1 API-Based Information Retrieval
-The Bing Visual Search API is used for the task of obtaining web pages related to a given image (vis, ) .
-A given image from the dataset is used to obtain a list of web pages completely and partially related to the image.
-We take the top three matching web pages in which the image appears.
-We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used.
-Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al., 2021), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results.
-In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base.
-Since this is relevant for only a very small number of examples in the dataset, it does not adversely affect system performance.
-3.3.2 Summarization using LLM
-Once the top three web pages are identified, we scrape their text to gather contextual information about the image.
-This textual data is often too lengthy for direct input to the agent, so we use the Llama-13B (Touvron et al., 2023) model to generate concise summaries, focusing on key details to help the agent better understand the external context.
-For non-English web pages, the LLM struggles with summarization, so we add a filter to exclude them.
-Although this restricts the system to English, it does not impact performance due to the dataset’s predominantly English-language content.
-Multi-lingual support could be added by translating text to English before summarization.
-3.4 Coherent Reasoning
-All the different components of LLM-Consensus are brought together in this stage of the pipeline.
-Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same.
-The agents also have access to external information related to the image through the external retrieval module.
-The final decision of the system is obtained once the debate terminates, which is after a certain number of debate rounds or after all agents converge to a common response, whichever is earlier.
+A model’s world knowledge is bounded by its training data and time frame, but external retrieval enables access to up-to-date context.
+Prior work leverages retrieval-based datasets [1], though these typically include only news headlines, which are often too sparse for reliable inference.
+Full article access can significantly improve misinformation detection.
+To this end, we introduce a two-stage external retrieval module that enhances accuracy when integrated into the pipeline:
+API-Based Information Retrieval
+We use the Bing Visual Search API to retrieve web pages related to each image.
+For each image, we collect the top three matching pages in which it appears, assuming these provide sufficient context for understanding the image’s original use.
+Since NewsCLIPpings [25] includes articles over a decade old, some images yield no search results.
+In these rare cases, we omit external context and rely solely on the agent’s prior knowledge, which has minimal impact on overall performance.
+Summarization using LLM
+After retrieving the top three web pages, we scrape their text and use LLaMA-13B [39] to generate concise summaries focused on key contextual details.
+Since the model struggles with non-English content, we filter out non-English pages–limiting the system to English but without affecting performance, as the dataset is predominantly English.
+Multilingual support could be added via translation prior to summarization.
+Coherent Reasoning
+This stage of the pipeline integrates all components of MAD-Sherlock.
+Each multimodal agent engages in the optimal debate setup using relevant prompts and is tasked with determining and explaining whether a given image-text pair constitutes misinformation.
+Agents also access contextual information via the external retrieval module.
+The system’s final decision is produced when the debate concludes, which is either after a fixed number of rounds or once all agents converge on a shared response, whichever occurs first.
 4 Experiments and Results
 4.1 Dataset
-We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al., 2021).
-The dataset is built based on the VisualNews (Liu et al., 2020) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian, and The Washington Post.
-The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair.
-CLIP (Radford et al., 2021) is used to retrieve semantically similar images for a given caption.
-We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples.
-The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively.
-We use the NewsCLIPpings dataset for evaluating our model’s efficacy in identifying out-of-context images and information due to its well-established use in prior work, its relative scale compared to other available datasets, and the lack of more recent large-scale benchmarks in this domain.
-Though no single dataset can capture the full diversity of real-world misinformation, NewsCLIPpings provides a standardized evaluation setting that facilitates comparability with existing methods.
+We conduct experiments on three datasets: NewsCLIPpings, VERITE, and MMFakeBench [25, 32, 24].
+This selection balances well-established benchmarks like NewsCLIPpings with newer, more diverse datasets such as VERITE and MMFakeBench, providing both continuity with prior work and coverage of recent advances.
+The NewsCLIPpings dataset builds on VisualNews [21], which contains image-caption pairs from BBC, USA Today, The Guardian, and The Washington Post.
+OOC samples are created by replacing an image with a semantically related one from a different pair.
+We use the Merged-Balanced version, with 71,072 training, 7,024 validation, and 7,264 test samples, offering balanced retrieval strategies and label distributions.
+Its scale and prior adoption make it well-suited for evaluating our model’s ability to detect out-of-context content.
+We also evaluate on the VERITE dataset [32], a recent benchmark for multimodal misinformation detection that controls for unimodal bias.
+Each image-caption pair is constructed so that neither modality alone determines veracity.
+VERITE includes three subsets—All, True vs.
+OOC, and True vs.
+MC—each testing different aspects of multimodal reasoning.
+Captions and images appear in both truthful and misleading contexts, promoting balanced, joint reasoning.
+The test split contains 338 true, 324 OOC, and 338 MC samples.
+We report results on the combined misinformation set (MC + OOC), with disaggregated results in Appendix A.4.
+We also evaluate on MMFakeBench [24], a benchmark for multimodal misinformation detection in mixed-source scenarios.
+The validation set includes 1,000 image-caption pairs spanning three major categories—textual, visual, and cross-modal distortions—each with 12 subtypes, capturing the complexity of real-world misinformation.
 4.2 Experimental Setup
-All experiments were run on 8 A40 (46GB) Nvidia GPU server. The estimated cost of processing one data sample using LLM-Consensus is \$0.24. Inference times range from 5 to 15 seconds.
-Debate Setup:
-We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al., 2024b).
-The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset.
-All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
-External Retrieval Module:
-We use the Bing Visual Search API (vis, ) to run an image-based reverse search.
-Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new, ).
-Finally, we use Llama-13B (Touvron et al., 2023) to summarise the text obtained from the top $k=3$ web pages.
+All experiments were run on 8 A40 (46GB) Nvidia GPU server.
+The estimated cost of processing one data sample using MAD-Sherlock with a GPT-4o backbone is \$0.24.
+Inference times range from 5 to 15 seconds.
+Debate Setup
+We conduct experiments to select the best debating configuration using the LLaVA model [23]. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
+External Retrieval Module
+We use the Bing Visual Search API to run an image-based reverse search.
+Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library.
+Finally, we use Llama-13B [39] to summarise the text obtained from the top $k=3$ web pages.
 This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
-Baselines and Prior Work:
-We compare LLM-Consensus to existing pretrained multi-modal baselines including CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2024b).
-We also compare performance against GPT-4o (OpenAI & et al., 2024; OpenAI, ).
+Baselines and Prior Work
 The models are presented with the image and caption pair and asked if the pair is misinformation.
 The models are further prompted to explain their reasoning.
-We also show results for two baseline methods trained from scratch, namely EANN (Wang et al., 2018) and SAFE (Zhou et al., 2020).
-We further compare LLM-Consensus to DT-Transformer (Papadopoulos et al., 2023), CCN (Abdelnabi et al., 2022), Sniffer (Qi et al., 2024), VINVL (Huang et al., 2024), SSDL (Mu et al., 2023) and Neuro-Sym (Zhu et al., 2022).
+We present comparisons to other explainable methods, specifically Sniffer [33], CRAVE [10], and MMD-Agent [24] in the following section.
+We also compare MAD-Sherlock to existing pretrained multi-modal baselines including CLIP [34], VisualBERT [19], InstructBLIP [9], LLaVA [23], and GPT-4o [30, 29] on the NewsCLIPpings dataset in Appendix A.4.
+We show results for two baseline methods trained from scratch on the NewsCLIPpings, namely EANN [41] and SAFE [44], in Appendix A.4.
+We further compare MAD-Sherlock to DT-Transformer [31], CCN [1], VINVL [17], SSDL [28], and Neuro-Sym [45] in Appendix A.4.
 4.3 Results
-We present results for the experiments conducted to select the best debate setup as well compare the performance of LLM-Consensus against existing methods.
+We present results for the experiments conducted to select the best debate setup as well compare the performance of MAD-Sherlock against existing methods.
 We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
 4.3.1 Comparing Debate Setups
-We compare multiple debating setups using the LLaVA model, to select the best one for comparison with other works and further experimentation.
+We compare multiple debating setups using the LLaVA-v1.6-34B model, to select the best one for comparison with other works and further experimentation.
 Debate Setup Accuracy Precision Recall
+Actor-Skeptic 69.5 66.1 69.4
+Judged Debate 66.7 66.7 61.5
 Async_DebateAI (believes debating AI) 75.2 54.5 86.4
@@
 Async_Debatehuman(w external info) 86.2 82.6 90.6
-Actor-Skeptic 69.5 66.1 69.4
-Judged Debate 66.7 66.7 61.5
-Debate with Disambiguation 77.8 74.7 82.6
-Table 1: Performance comparison between different debate setups: The Async_Debatehuman where the model has external context and believes it is debating a human being is the best setup.
-Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations.
-We also report results for the Async_Debatehuman setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain.
-The external retrieval of information significantly boosts performance.
-We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent.
-Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human.
-Further, the asynchronous debate setup benefits from the ensemble of agents, which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses.
-The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought.
-We believe this accounts for the counter-intuitively performance of this method where agents perform worse with more information.
-The judged debate setup focuses on enforcing agents to structure their responses in a way that will convince a judge.
-The agents also debate with opposite stances and do not have the option of changing their stance mid-debate.
-This can further confuse the judge and lead to incorrect decisions.
-This is resolved in the Async_Debatehuman set up where agents are given complete freedom over their initial opinions, as well as their opinions during the debate.
-If they believe they are convinced by the other agents’ arguments, they can choose to change their response and the debate ends.
-Based on the results from Table 1, we choose the best-performing debate set-up, i.e.
-Async_Debatehuman (with external information) as the debate configuration for further experimentation and comparison.
+Table 1: Performance comparison between different debate setups: We observe best performance with the The Async_Debatehuman set-up, in which the model has external context and believes it is debating a human.
+Table 1 shows that Async_Debatehuman with external information outperforms all other debate configurations.
+To highlight the role of external retrieval, we also report results without it and observe a significant performance drop.
+We find that agents perform better when they believe they are debating a human, engaging more critically with peer responses.
+The asynchronous setup also benefits from ensemble effects, unlike the actor-skeptic setup where only one agent generates responses.
+In judged debates, agents must adhere to fixed stances and persuade a judge, which can lead to confusion and errors.
+In contrast, Async_Debatehuman allows agents to revise their views mid-debate, leading to clearer outcomes.
+Based on these findings, we adopt this setup—with external information—for all further experiments.
 4.3.2 Performance Comparison
-We present our results on the NewsCLIPpings dataset against existing out-of-context detection methods discussed in section 4.2.
+We present our results on the NewsCLIPpings, VERITE, and MMFakeBench datasets against existing out-of-context detection methods discussed in Section 4.2.
+We include additional comparisons against legacy methods in Appendix A.4.
+Dataset Model Performance
+Accuracy Precision Recall
+NewsCLIPpings SNIFFER [33] 88.4 91.8 86.9
+CRAVE [10] 85.0 85.0 85.0
+MAD-Sherlock (ours) 90.8 85.5 99.0
+VERITE CRAVE [10] 82.0 82.0 82.0
+MAD-Sherlock (ours) 85.2 83.6 96.0
+MMFakeBench MMD-Agent [24] 62.1 67.8 59.3
+CRAVE [10] 78.0 83.6 69.7
+MAD-Sherlock (ours) 83.8 87.2 90.1
+Table 2: Performance comparison of LLM-Consensus against other explainable methods on various datasets: Performance of LLM-Consensus and baselines across NewsCLIPpings, VERITE, and MMFakeBench.
+Our model shows consistent generalization and superior performance across datasets.
+Table 2 shows the comparison between our method and other explainable methods on the NewsCLIPpings, VERITE, and MMFakeBench datasets.
+We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o [30, 29] model across all datasets.
+As a reminder, Sniffer [33] is finetuned extensively to adapt it to the NewsCLIPpings dataset.
+MMD-Agent represents the single-agent framework proposed by the authors of the MMFakeBench dataset [24].
+The CRAVE framework introduced by Dey et al.
+[10] represents the most recent work across all three datasets, but our framework shows markedly superior performance across all datasets.
+We find that our system produces coherent, detailed and comprehensive explanations.
+We also note that the debate paradigm in itself is essential to the system performance.
+We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model.
+We believe that this follows from work by Mireshghallah et al.
+[27], which introduces and demonstrates the importance of contextual privacy.
+Introducing multiple context windows allows each agent to maintain its own role and informational perspective without leakage, which is not possible in single-agent systems.
+We provide a qualitative sample of our framework’s explanations in Appendix A.5. We provide ablation studies on each component of the MAD-Sherlock pipeline in Appendix A.6.
+We also note that single multi-modal models that do not do retrieval, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work.
+We demonstrate these results in Appendix A.4.
+This can be attributed to the necessity for external context for misinformation detection and the lack of diverse perspectives that occur naturally in a multi-agent framework.
+These models require additional integration into more comprehensive pipelines, as done in this work.
+5 User Study
+We conducted a user study to assess our system’s ability to detect and explain misinformation—particularly important given the lack of standard metrics for evaluating explanation quality.
+Participants were grouped by profession: Journalists, Academics, and Others (see Appendix A.7).
+Each participant reviewed ten image-text pairs, judged their veracity, and rated their confidence (0–10).
+After submitting initial responses, they viewed MAD-Sherlock’s explanations and revised their answers.
+As shown in Table 3, MAD-Sherlock outperformed average human accuracy both with and without AI assistance, highlighting its potential for improving public safety and trust.
+Study Setup Average Accuracy
+Humans $60.3\pm 13.5$
+Humans+MAD-Sherlock $76.7\pm 12.2$
+MAD-Sherlock $\mathbf{80.0\pm 0.0}$
+Table 3: Performance comparison between different study setups: MAD-Sherlock outperforms humans with and without AI assistance.
+Group-wise analysis, shown in Table 4, reveals significant performance gains across all groups, with results approaching those of professional journalists.
+Confidence levels (out of 10) are comparable across groups and generally increase after using MAD-Sherlock insights.
+Thus, MAD-Sherlock can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
+Metric Journalists Academics Others
+Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
+Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
+Accuracy (with MAD-Sherlock) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
+Confidence (with MAD-Sherlock) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
+Table 4: Performance comparison: MAD-Sherlock improves performance across all participant groups.
+6 Conclusion and Future Work
+Out-of-context (OOC) image misuse is an increasing challenge for misinformation detection, especially as vision-language models grow more powerful yet less interpretable.
+We explore whether multiple AI agents can collaboratively reason about context to improve prediction accuracy.
+Our strongest results come from the Asynchronous_Debatehuman setup, where agents believe they are debating a human.
+This setting promotes engagement, mid-debate revision, and better identification of subtle inconsistencies.
+Our final system, MAD-Sherlock, achieves state-of-the-art performance while offering interpretable, evidence-based explanations—enabled by our external retrieval module.
+We observe substantial gains in OOC detection across both expert and non-expert users.
+We see several avenues for future work, including a continuously updated benchmark with recent news and more nuanced inconsistencies, extending to video-text and multilingual inputs, and large-scale deployments in professional and citizen intelligence settings.
+We hope to engage with the community to better understand how agentic workflows can enhance online safety.
+For limitations, see Appendix A.1.
+Appendix A Appendix
+A.1 Limitations
+Despite the strong performance of MAD-Sherlock, several limitations remain.
+First, while our model excels at detecting out-of-context image–text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve.
+Moreover, our framework cannot independently verify the factual correctness of externally retrieved news articles; the truthfulness of any source may be debated, potentially introducing misinformation into the reasoning process.
+Nevertheless, we observe that the model’s judgments align closely with the human-created labels used in these widely accepted datasets, underscoring its practical utility.
+Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration.
+Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
+Additionally, while our user studies provided valuable insights, large-scale deployment in diverse, real-world settings, such as professional or citizen intelligence environments, is necessary to fully assess the method’s robustness and scalability.
+Finally, our dataset, though comprehensive, primarily focuses on English-language news, limiting the generalizability of the system across non-English contexts.
+Another important limitation is the potential risk that open-sourcing MAD-Sherlock might allow adversaries to train models specifically designed to counter or evade detection by our system.
+As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness.
+However, despite these risks, we believe that open-sourcing remains the right path forward.
+Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
+Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining MAD-Sherlock’s effectiveness in the long term.
+The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
+Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of MAD-Sherlock.
+A.2 Sample Image-Caption Pair in the News Domain
+Figure 4: Russian President Vladimir Putin has called Ukraine’s move into Kursk a “major provocation”.
+Image and caption taken from the BBC article here (Accessed at 17:43 on Aug 11, 2024): https://www.bbc.co.uk/news/articles/cze5pkg5jwlo
+A.3 Prompts for MAD-Sherlock
+You are a misinformation detection expert in the news domain.
+You will look at image-caption pairs and decide if the given image is rightly used in the given news context.
+To further assist you, a summary of news articles related to the image will be provided.
+Based on this, you need to decide if the caption belongs to the image or if it is being used to spread false information to mislead people.
+Note that the image is real.
+It has not been digitally altered.
+Carefully examine the image for any known entities, people, watermarks, dates, landmarks, flags, text, logos and other details which could give you important information to better explain your answer.
+Remember in news articles images and captions are often related contextually and the caption need not exactly describe the image.
+The goal is to consider the contextual relationship between the image and caption based on the news articles and correctly identify if the image caption pair is misinformation or not and to explain your answer in detail.
+Think step by step and plan a detailed explanation for your answer.
+Figure 5: Model initialization prompt.
+This is a summary of news articles related to the image: {}
+Based on this, you need to decide if the caption given below belongs to the image or if it is being used to spread false information to mislead people.
+CAPTION: {}
+Note that the image is real. It has not been digitally altered.
+Carefully examine the image for any known entities, people, watermarks, dates, landmarks, flags, text, logos and other details which could give you important information to better explain your answer.
+The goal is to correctly identify if this image caption pair is misinformation or not and to explain your answer in detail.
+At the end give a definite YES or NO answer to this question:
+IS THIS MISINFORMATION?
+Figure 6: Initial prompt for independent opinion formation and response generation with GPT-4o.
+This is what I think: {}.
+Do you agree with me?
+If you think I am wrong then convince me why you are correct.
+Clearly state your reasoning and tell me if I am missing out on some important information or am making some logical error.
+Do not describe the image.
+At the end give a definite YES or NO answer to this question:
+IS THIS MISINFORMATION?
+Figure 7: Prompt for first round of debate.
+I see what you mean and this is what I think: {}.
+Do you agree with me?
+If not then point out the inconsistencies in my argument (e.g. location, time or person related logical confusion) and explain why you are correct.
+If you disagree with me then clearly state why and what information I am overlooking.
+Find disambiguation in my answer if any and ask questions to resolve them.
+I want you to help me improve my argument and explanation.
+Don’t give up your original opinion without clear reasons, DO NOT simply agree with me without proper reasoning.
+At the end give a definite YES or NO answer to this question:
+IS THIS MISINFORMATION?
+Figure 8: Prompt for subsequent rounds of debate.
+A.4 Supplemental Results
+A.4.1 Additional Results on VERITE
+Model External Retrieval Data Accuracy Precision Recall
+4o VERITE (all) 85.2 83.6 96.0
+4o VERITE (true vs. OOC) 79.5 69.7 96.4
+4o VERITE (true vs. MC) 80.7 73.6 95.6
+4o VERITE (all) 84.8 85.0 93.1
+4o VERITE (true vs. OOC) 78.6 72.1 89.5
+4o VERITE (true vs. MC) 82.6 75.6 96.4
+Table 5: External Retrieval Ablation on VERITE.
+We compare performance of LLM-Consensus with and without external retrieval with a GPT-4o backbone.
+We also disaggregate our results by “miscaptioned“ (MS) and “out-of-context“ (OOC), two different classes of misinformation included in the VERITE dataset.
+A.4.2 Additional Baselines for NewsCLIPings
 Model Accuracy $\uparrow$
@@
 Sniffer (w/o finetuning) 84.5
-LLM-Consensus (ours) 90.8
-Table 2: Performance comparison between our model and baselines: LLM-Consensus (with GPT-4o) out performs all related work.
-Note: the GPT-4o# setup is identical to LLM-Consensus with the absence of a multi-agent debate, here only a single agent with access to external information is considered.
-Table 2 shows the comparison between our system and existing methods.
-We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al., 2024; OpenAI, ) model.
-Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset.
-While we do not provide a quantitative assessment of explanations by LLM-Consensus, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines.
-This is attributed to the fact that in a multi-agent setup, we have multiple context windows which leads to more coherent and relevant final explanations.
-We leave the detailed analysis of these explanations and the development of the associated metrics as future work.
-We also note that the debate paradigm in itself is essential to the system performance.
-We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model.
-We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work.
-This can be attributed to the necessity for external context for misinformation detection in the news domain and the lack of diverse perspectives that arise naturally in a multi-agent framework.
-Therefore, these standalone models, while promising, currently are unable to detect misinformation effectively.
-These models require additional integration into more comprehensive pipelines, as done in this work.
-In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al., 2020) and EANN (Wang et al., 2018) perform worse than pretrained multi-modal models.
-This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
-4.3.3 Ablations
-To analyze the importance of each component of the LLM-Consensus framework, we conduct ablation experiments.
+MAD-Sherlock (ours) 90.8
+Table 6: Performance comparison on NewsCLIPpings against various legacy methods: MAD-Sherlock (with GPT-4o) outperforms all related work.
+Note: the GPT-4o# setup is an ablation in which there is a single agent using our external retrieval module (no debate).
+A.5 Qualitative Analysis
+Image Description: A banner that says ‘‘STOP VICTIMISATION OF UNION MEMBERS END THE BLACKLIST’’ in front of a city landscape.
+Caption: Flashmob protest at Crossrail site in Central London by Unite union members in protest against blacklisting and sackings of union officials and members.
+Label: This is not misinformation.
+Framework prediction: This is not misinformation.
+Explanation (Model Output): Here is a precise breakdown:
+1. Image Analysis:
+- The banner in the image clearly states ‘‘STOP VICTIMISATION OF UNION MEMBERS END THE BLACKLIST.’’
+- This message aligns with union-related grievances focused on stopping blacklisting and mistreatment of union members.
+2. Caption Analysis:
+- The caption explicitly mentions a flashmob protest by Unite union members at a Crossrail site in Central London.
+- The topics of blacklisting and the sacking of union officials and members are mentioned in the caption, aligning with the grievances captured in the image.
+3. News Summary Correlation:
+- The true news articles discuss blacklisting practices within the construction industry, particularly concerning the Crossrail project and actions taken by the Consulting Association.
+- Gail Cartmail from Unite the Union has provided evidence of blacklisting affecting union members working on the Crossrail project, which fits the context depicted in the banner.
+- The articles also highlight broader issues of blacklisting, further supporting the theme reflected in the image and caption.
+4. Contextual Match:
+- The image and the caption are aligned in context. The message on the banner directly supports the protest described in the caption.
+- The news article summary corroborates the issues of blacklisting faced by union members, as illustrated by the protest.
+Your reasoning is logically sound, and no critical information is missing.
+Both the image and the caption are contextually connected to the theme of union protests against blacklisting, specifically within the Crossrail project setting.
+Thus, I conclude:
+Definite Answer: NO, the image-caption pair does not represent misinformation.
+Figure 9: Qualitative sample of model explanation.
+A.6 Ablations
+A.6.1 Ablations
+To analyze the importance of each component of the MAD-Sherlock framework, we conduct ablation experiments.
 Specifically, we evaluate the effect of using LLaVA against GPT-4o, the impact of the external retrieval module, and the power of the multi-agent debate framework.
@@
 $\mathbf{90.8}$ $85.5$ $\mathbf{99.0}$
-Table 3: Ablation: Quantitative evaluation of each component of LLM-Consensus on classification performance.
+Table 7: Ablation: Quantitative evaluation of each component of MAD-Sherlock on NewsCLIPpings classification performance.
 We observe that the combination of GPT-4o, the external retrieval module, and the multi-agent debate framework yields the highest performance across all metrics, with 90.8% accuracy, 85.5% precision, and 99.0% recall, demonstrating the value of combining these components.
-The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7 (as seen in Table 2) to 90.2, underscoring its role in enabling contextual reasoning and refining predictions.
-Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3 to 85.5 and recall from 90.1 to 99.0.
+The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7% (as seen in Table 6) to 90.2%, underscoring its role in enabling contextual reasoning and refining predictions.
+Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3% to 85.5% and recall from 90.1% to 99.0%.
 Meanwhile, retrieval contributes more substantially to LLaVA’s performance gains, likely due to GPT-4o’s broader world knowledge.
 Without external retrieval or the debate framework, the performance drops, emphasizing the critical role of these components in achieving state-of-the-art results.
-5 User Study
-We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation.
-While model performance in misinformation detection is easy to quantify, there are no established metrics for assessing the quality of its explanations, making the user study essential.
-Participants were grouped into three categories based on their professions: Journalists, AI Academics, and Others (see Appendix A.4 for details).
-During the study, participants reviewed ten image-text pairs and decided whether each pair constituted misinformation.
-They also rated their confidence on a scale of 0 to 10.
-After submitting their initial answers, participants were shown LLM-Consensus’s AI-generated explanations and asked to reconsider their responses.
-As shown in Table 4, the system outperformed the average human performance in both cases, with and without AI insights.
-These results demonstrate LLM-Consensus’s potential as a reliable assistive tool for OSINT research, capable of detecting and explaining misinformation with minimal human intervention.
-[…]
-Humans $60.3\pm 13.5$
-Humans+LLM-Consensus $76.7\pm 12.2$
-LLM-Consensus $\mathbf{80.0\pm 0.0}$
-Table 4: Performance comparison between different study setups: LLM-Consensus outperforms humans with and without AI assistance.
-Group-wise analysis, shown in Table 5, reveals significant performance gains across all groups, with results approaching those of professional journalists.
-Confidence levels (out of 10) are comparable across groups and generally increase after using LLM-Consensus insights.
-Thus, LLM-Consensus can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
-Metric Journalists Academics Others
-Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
-Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
-Accuracy (with LLM-Consensus) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
-Confidence (with LLM-Consensus) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
-Table 5: Performance comparison: LLM-Consensus improves performance across all participant groups.
-6 Conclusion and Future Work
-Misinformation detection has become a pressing issue in recent times.
-With the ever-advancing capabilities of vision and language models, the detection of OOC image use has become a very difficult task.
-In this work, we explore the question of whether it is possible for multiple AI agents to pool their contextual knowledge and converge to a common prediction in order to identify instances of misinformation.
-We identify Asynchronous_Debatehuman as the most optimal communication setup for AI models.
-We observe significant performance improvement when the models believe they are debating against a human instead of another AI agent.
-We observe that in this setup, models tend to be more involved and open to changing their opinions.
-Our method also allows for agents to have freedom of opinion which they may change mid-debate.
-Agents in such a setting show enhanced abilities to critically evaluate an argument and pick up on minute inconsistencies.
-Our final system, LLM-Consensus, achieves state-of-the-art performance in misinformation detection while offering clear, detailed explanations, thanks to our advanced external retrieval module.
-As a result, we observe significant improvements in out-of-context (OOC) detection for both experts and non-experts.
-We identify several promising directions for future research.
-The community would benefit from a continuously updated benchmark dataset incorporating recent news articles and subtler inconsistencies.
-We plan to extend our methods to video-text pairs and multi-lingual content.
-We also believe that integrating more advanced models in the summarization pipeline will further enhance performance and are interested in comparing LLM-Consensus with multi-agent systems using external retrieval.
-Finally, we conducted extensive user studies but recognize the importance of large-scale deployment in professional settings and the citizen intelligence community.
-This will provide valuable real-world feedback and uncover further opportunities for improvement.
-For a discussion of limitations, see Appendix A.1.
-Appendix A Appendix
-A.1 Limitations
-Despite the strong performance of LLM-Consensus, several limitations remain.
-First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve.
-Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration.
-Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
-Additionally, while our user studies provided valuable insights, large-scale deployment in diverse, real-world settings, such as professional or citizen intelligence environments, is necessary to fully assess the method’s robustness and scalability.
-Finally, our dataset, though comprehensive, primarily focuses on English-language news, limiting the generalizability of the system across non-English contexts.
-Another important limitation is the potential risk that open-sourcing LLM-Consensus might allow adversaries to train models specifically designed to counter or evade detection by our system.
-As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness.
-However, despite these risks, we believe that open-sourcing remains the right path forward.
-Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
-Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining LLM-Consensus’s effectiveness in the long term.
-The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
-Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of LLM-Consensus.
-A.2 Sample Image-Caption Pair in the News Domain
-Figure 4: Russian President Vladimir Putin has called Ukraine’s move into Kursk a “major provocation”.
-Image and caption taken from the BBC article here (Accessed at 17:43 on Aug 11, 2024): https://www.bbc.co.uk/news/articles/cze5pkg5jwlo
-A.3 Prompts for LLM-Consensus
-This is a summary of news articles related to the image: {}
-Based on this, you need to decide if the caption given below belongs to the image or if it is being used to spread false information to mislead people.
-CAPTION: {}
-Note that the image is real. It has not been digitally altered.
-Carefully examine the image for any known entities, people, watermarks, dates, landmarks, flags, text, logos and other details which could give you important information to better explain your answer.
-The goal is to correctly identify if this image caption pair is misinformation or not and to explain your answer in detail.
-At the end give a definite YES or NO answer to this question:
-IS THIS MISINFORMATION?
-Figure 5: Initial prompt for independent opinion formation and response generation
-This is what I think: {}.
-Do you agree with me?
-If you think I am wrong then convince me why you are correct.
-Clearly state your reasoning and tell me if I am missing out on some important information or am making some logical error.
-Do not describe the image.
-At the end give a definite YES or NO answer to this question:
-IS THIS MISINFORMATION?
-Figure 6: Prompt for Debate Round 1
-I see what you mean and this is what I think: {}.
-Do you agree with me?
-If not then point out the inconsistencies in my argument (e.g. location, time or person related logical confusion) and explain why you are correct.
-If you disagree with me then clearly state why and what information I am overlooking.
-Find disambiguation in my answer if any and ask questions to resolve them.
-I want you to help me improve my argument and explanation.
-Don’t give up your original opinion without clear reasons, DO NOT simply agree with me without proper reasoning.
-At the end give a definite YES or NO answer to this question:
-IS THIS MISINFORMATION?
-Figure 7: Prompt for Debate after Round 1
-A.4 User Study
+We show additional results on VERITE when ablating the retrieval module in Appendix A.4.
+A.7 User Study
 We conduct a user study to assess the effectiveness of our model in detecting and explaining misinformation. Through this study, we aim to assess the persuasiveness of our system.
-A.4.1 Setup
+A.7.1 Study Design
 The user study was designed to evaluate the effectiveness of our system in detecting and explaining misinformation.
@@
 A total of 30 participants volunteered to participate in this study.
-The group of individuals included journalists from BBC as well as students and professors from the University of Oxford.
 Participation was completely voluntary and no personal information was used for the purpose of analysis in this study.
@@
 For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair.
-These AI insights were the final outputs from LLM-Consensus.
+These AI insights were the final outputs from MAD-Sherlock.
 Participants were then asked to reconsider their answer and again decide if the image-text pair was misinformation or not, in light of the new information from the AI agent.
@@
 This was done to ensure an unbiased estimate of average human performance.
-The image-text pairs to include in the study were taken from the NewsCLIPpings (Luo et al., 2021) dataset.
+The image-text pairs to include in the study were taken from the NewsCLIPpings [25] dataset.
 AI insights were taken from our best-performing setup involving the GPT-4o model.
@@
 Therefore the model accuracy for the task was 80% and we use this as the baseline accuracy to compare human performance against.
-We analyse two special cases, where LLM-Consensus argues for the wrong answer.
+We analyse two special cases, where MAD-Sherlock argues for the wrong answer.
 We include these results in order to observe how persuasive our system can be even when it is wrong.
-We note in the instance where the image-text pair was actually misinformation and the model argued that it was not, 6 participants changed their correct responses to those suggested by LLM-Consensus.
+We note in the instance where the image-text pair was actually misinformation and the model argued that it was not, 6 participants changed their correct responses to those suggested by MAD-Sherlock.
 Although this is only 5% of the participants, it still gives a significant insight into how persuasive the model can appear even when it is wrong.
 While the case of false negatives is important, false positives are an even more concerning matter for our problem statement.
-In the case where LLM-Consensus declared the given image-text pair to be misinformation when it was not, is important to analyse.
+In the case where MAD-Sherlock declared the given image-text pair to be misinformation when it was not, is important to analyse.
 In this setting 50% of the total participants changed their answer to the wrong one, therefore believing a piece of true information to be false.
@@
 The same task without AI insights would require extensive analysis and we project it would take between 30-45 minutes to decide if 10 image-text pairs were misinformation.
-A.5 Multi-modal debates for hamful meme detection
+A.8 Screenshots
+We include representative screenshots of the user study.
+(a) Instructions screen
+(b) No internet baseline
+(c) With internet baseline
+(d) With retrieval-augmented summary
+Figure 10: Screenshots of the user survey. The questions asked after AI summary is presented are the same as those following the other questions.
+A.9 Multi-modal debates for harmful meme detection
 While this work relates to a different problem than OOC misinformation detection in the news domain, we still find the approach taken by the authors a relevant related work and therefore include it here.
-(Lin et al., 2024) use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful.
+Lin et al.
+[20] use LMMs debating against each other to generate explanations for contradictory arguments regarding whether a given meme is harmful.
 These explanations are then used to train a small language model as a judge to determine whether the image and text that make up the meme are actually harmful.
@@
 Finally, this work is related to harmful meme detection and does not concern the problem of misinformation detection in the news domain, which likely requires more intricate contextual analysis, including of external context.
+A.10 Additional Experiments
+Debate with Disambiguation:
+Building on the actor-skeptic method, we allow all agents to act as both actors and skeptics.
+Models generate their own responses and disambiguation queries to refine or challenge other agents’ outputs.
+These queries are used to retrieve additional information from the Internet, further improving model responses.
+The Debate with Disambiguation strategy achieves accuracy of 77.8, precision of 74.7, and recall of 82.6 when tested with a LLaVa backbone.